### v1

Abstract
We find that the spontaneous and collective emissions have a large influence on the excitation of two-level absorbers (atoms, molecules) interacting in resonance with the plasmonic mode near the metal nanoparticle. The spontaneous and collective emissions restrict the absorption enhancement by plasmonic mode and make the enhancement possible only with a fast, on a pico-second time scale, population relaxation of the upper absorbing states. Conditions for the maximum of the plasmon-enhanced absorption in the presence of spontaneous and collective emissions are found. We show that the nanoparticle-absorber interaction and the collective emission cause the bistability in the plasmon-enhanced absorption.
I Introduction
The localized plasmon resonances (LPR) excitation leads to the electric field enhancement near metal nanoparticles Maier (2007); Pelton, Aizpurua, and Bryant (2008); Amendola et al. (2017); Tanabe (2008). Such a "plasmonic" field enhancement finds various applications in optical sensors Li, Cushing, and Wu (2015), bio-medicine Liao, Nehl, and Hafner (2006), photovoltaics Jang et al. (2016), photo-catalysis Zhang et al. (2013) and other areas of physics and chemistry Maier (2007); Barbillon (2019). A new branch of chemistry - plasmonic chemistry was developed concerning the chemical reactions Zhan et al. (2023); Cortés (2017) stimulated by the electric field enhanced near plasmonic nanoparticles.
A photochemical reaction is initiated by the absorption of light and the creation of transient molecules in excited states whose properties differ greatly from the original molecules. These new chemical species can fall apart, change to new structures, combine other molecules, or transfer electrons, hydrogen atoms, protons, or their electronic excitation energy to other molecules Fleming, Krueger, and Longworth (2023). So the optimization of the plasmonic photo-chemical reactions is important with the searching for conditions of the maximum field and absorption enhancement near metal nanoparticles Kelly et al. (2003). This paper contributes to such an optimization.
A large electric field enhancement is well-known near the sharp features of nanostructures, as "hot spots" Díaz-Núñez et al. (2019), or in a small gap between nanoparticles Kim et al. (2010). However, it may be difficult to make a sufficiently large number of "hot spots" or small gaps in nanostructures. So in this paper, we do not consider the field enhancement by "hot spots" or in gaps between nanoparticles. We restrict ourselves by the absorption enhancement in the near field of metal nanoparticles with the excited LPR.
We will pay attention on the spontaneous and collective emissions, which significantly affect the resonant absorption near plasmonic nanoparticles and cause new physical effects. Spontaneous emission near metal nanoparticles has been investigated, for example, in Blanco and García de Abajo (2004). Collective effects as a super-radiance have been studied for the resonant absorbers, as atoms or molecules, in the near zone of the plasmonic particle Shahbazyan and Pustovit (2011); Protsenko and Uskov (2015); Protsenko et al. (2017). However, the absorption enhancement near the nanoparticle with an account for the spontaneous and collective emissions deserves more studies. The influence of spontaneous emission and collective effects on absorption is important for plasmonic chemistry, particularly, in the relation with recently studied plasmon-induced direct resonance energy transfer (PIRET) processes You, Ramakrishna, and Seideman (2018). A part of studies of the plasmon-enhanced absorption with the spontaneous and collective emissions to the LPR mode, with the direct energy transfer from the plasminic nanoparticle to the absorbing molecule, is done in the present paper.
Here we use the model for the resonant absorption near plasmonic nanoparticles similar to the laser model. The plasmonic nanoparticle makes a cavity for the electric field similar to the laser cavity, as in the plasmonic nanolaser Protsenko et al. (2005, 2008); Protsenko (2012). A part of our approach is similar to the analysis of the spontaneous emission and collective effects in lasers André et al. (2019); Protsenko and Uskov (2022).
The plasmonic mode provides good conditions for collective effects in the resonant absorption. Collective effects in the resonant medium are significant in the low-quality cavity André et al. (2019); Protsenko and Uskov (2022); Piccione, Felicetti, and Bellomo (2022). The LPR mode has a low quality, appropriate for collective effects. The particle surface is a good place for the resonant absorbers (atoms, molecules, q-dots, etc.).
In this paper, we show that the spontaneous emission into the LPR mode and collective interaction of absorbers through the plasmonic electric field significantly affect the absorber excitation in the vicinity of metal nanoparticles and lead to bistability in the excitation. We will find conditions for the best absorber excitation and bistability. These results are important for all branches of plasmonic chemistry, photovoltaic, and all applications related to the resonant excitation of absorbers near plasmonic nanoparticles.
We derive basic equations quantum-mechanically as Heisenberg operator equations. In the present paper, we solve these equations in the semi-classical approximation as it is common, for example, in the laser theory Sargent, Scully, and Lamb (1974). Fully quantum analysis of the equations will be done in the future.
We carry out calculations analytically. The analytical approach lets to find expressions illustrating the physical processes at the absorber excitation and the influence of the plasmonic particle on important parameters such as the polarization and the population of states of absorbers.
We consider the increase of the absorber excited state population as a criterion of absorption enhancement by the LPR.
In the next section, we describe the model, Hamiltonian, and equations of motion.
The expressions, characterizing the stationary absorption, as the absorption coefficients and the energy balance equation, are presented in section III. Several physical processes at the resonant absorption near the metal nanoparticle are explained with the help of equations found in section III.
Section IV shows conditions for the best plasmon-enhanced absorption and describes the bistability.
Examples of the absorption enhancement and the bistability are presented in section V at typical values of the nanoparticle and the absorber parameters.
Results are discussed in section VI, and the paper is finalized in the conclusion section VII.
II The model, Hamiltonian and equations of motion
II.1 The model
Suppose, a metal nanoparticle interacts with $i=1...N_{m}$ identical absorbers, as molecules or atoms, being on the distance $r$ from the particle center on the dielectric shell as shown in Fig. 1. The absorber transition frequency is resonant with the dipole LPR mode of the nanoparticle.
Figure 1: Coordinate system, the nanoparticle 1 and the resonant absorber $i$ (a molecule or an atom) in the external electric field $\hat{\vec{E}}$. $r$ is the distance, and $\vec{e}_{ri}$ is the unit vector from the nanoparticle center to the absorber. $\left|00\right>_{i}$ is the absorber ground state. The dipole momentum with the matrix element $d$ for the absorber transition from the excited state $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to $\left|00\right>_{i}$ is parallel to the unit vector $\vec{e}$ of the external field polarization (or to the vector $\vec{e}_{i}$). The absorber state with the transition dipole momentum parallel to the axis $OO^{\prime}$ is not excited; such transition does not interact with the external and the nanoparticle fields. The insert shows the resonant absorbers randomly distributed on the nanoparticle shell.
Such resonance can be achieved for the nanoparticle of appropriate shape Kelly et al. (2003). For simplicity, we consider a spherical nanoparticle.
The particle and absorbers interact with the external monochromatic electric field of the amplitude $\hat{\vec{E}}$ linearly polarized along the unit vector $\vec{e}$. The field $\hat{\vec{E}}$ is resonant with the LPR and the absorber transitions.
We consider the $i$-th absorber in the Cartesian coordinate system with the origin in the particle center. One coordinate axis is parallel to $\vec{e}$, another axis with the unit vector $\vec{e}_{i}$ is in the plane made by $\vec{e}$ and the unit vector $\vec{e}_{ri}$ directed from the nanoparticle center to the i-th absorber. The angle $\theta_{i}$ is between $\vec{e}$ and $\vec{e}_{i}$, and the third coordinate axis is along the line $OO^{\prime}$. Two excited ${{\left\langle 10\right|}_{i}}$, ${{\left\langle 01\right|}_{i}}$ and the ground ${{\left\langle 00\right|}_{i}}$ states of i-th absorber participate in the interaction with electric fields near the nanoparticle. Such absorber states are, for example, the lowest unoccupied molecular orbital (LUMO) and the highest occupied molecular orbital (HOMO) Fukui, Yonezawa, and Shingu (2004).
We will calculate the population of excited states of absorbers and examine how the interaction between the particle and absorbers affects the population.
We suppose a large number $N_{m}\gg 1$ of absorbers randomly distributed on the nanoparticle shell as shown on the insert in Fig. 1. In the case of $N_{m}\gg 1$, the absorber-absorber interaction through the near field leads to only the self-broadening of the absorber’s transitions Royer (1975). We include the self-broadening in the absorber transition width $\gamma$ and do not describe the absorber-absorber interaction explicitly.
II.2 Hamiltonian
Following Protsenko et al. (2005) we describe the nanoparticle as a dipole making harmonic oscillations in the resonant external electric field. We introduce the nanoparticle dipole momentum operator
$\hat{\vec{p}}=p\vec{e}\left(\hat{b}{{e}^{-i{{\omega}_{0}}t}}+{{{\hat{b}}}^{+}}{{e}^{i{{\omega}_{0}}t}}\right)$ (1)
where $p$ is the dipole momentum matrix element, $\hat{b}$ is a Bose-operator describing collective (plasmonic) oscillations of the nanoparticle electrons in the electric field. Later we express $p$ through the nanoparticle polarizability Protsenko et al. (2005). For symmetry reasons, we suppose the same directions for the nanoparticle dipole and the external field polarization.
The dipole momentum operator of the $i$-th absorber resonant transitions is
${{\hat{\vec{d}}}_{i}}=d\left[\left({{{\hat{\sigma}}}_{i}}\vec{e}+{{{\hat{\sigma}}}_{\bot i}}{{{\vec{e}}}_{i}}\right){{e}^{-i{{\omega}_{0}}t}}+h.c.\right].$ (2)
Here $d$ is the dipole transition matrix element; absorber operators in Dirac notations ${{\hat{\sigma}}_{i}}={{\left|00\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$, (or ${{{\hat{\sigma}}}_{\bot i}}={{\left|00\right\rangle}_{i}}{{\left\langle 10\right|}_{i}}$) are transition operators from the high $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to the low $\left|00\right>_{i}$ states of the absorber $i$ parallel to $\vec{e}$, (or $\vec{e}_{i}$): ${{{\hat{\sigma}}}_{i}}\left|01\right>_{i}=\left|00\right>_{i}$ and ${{{\hat{\sigma}}}_{\perp i}}\left|10\right>_{i}=\left|00\right>_{i}$.
Hamiltonian of the particle and $N_{m}$ absorbers in the electric field $\hat{\vec{E}}$ is
$H={{H}_{0}}-\left(\hat{\vec{p}}+\sum\limits_{i=1}^{N}{{{{\hat{\vec{d}}}}_{i}}}\right)\hat{\vec{E}}+{{\hat{V}}_{int}}+\hat{\Gamma}$ (3)
The first term in (3) describes free oscillations of dipoles; the second term is for the dipole interaction of the particle and absorbers with the external field; the third term ${{\hat{V}}_{int}}$ describes the interaction of the particle and absorbers through the nanoparticle near field (the LPR field) and the last term in (3) describes the dissipation and the interaction of the particle and absorbers with baths.
We suppose that the distance $r$ between absorbers and the particle center is much smaller than the LPR wavelength. So the particle-absorber interaction ${{\hat{V}}_{int}}$ is a quasi-static dipole-dipole interaction. We take the dipole-dipole quasi-static interaction energy from Landau and Lifshitz (1982), replace classical dipoles with operators, and write
${{\hat{V}}_{int}}=\frac{1}{{{r}^{3}}}\sum\limits_{i=1}^{{{N}_{m}}}{\left[\hat{\vec{p}}{{{\hat{\vec{d}}}}_{i}}-3(\hat{\vec{p}}{{{\vec{e}}}_{ri}})({{{\hat{\vec{d}}}}_{i}}{{{\vec{e}}}_{ri}})\right]},$ (4)
where ${{{\vec{e}}}_{ri}}$ is the unit vector along the direction from the nanoparticle center to i-th absorber as shown in Fig. 1.
The external field is
$\hat{\vec{E}}={{E}_{0}}\vec{e}\left(\hat{a}{{e}^{-i{{\omega}_{0}t}}}+{{{\hat{a}}}^{+}}{{e}^{i{{\omega}_{0}}t}}\right)$ (5)
where $\hat{a}$ is the field amplitude Bose-operator, ${{E}_{0}}=\eta^{-1}\sqrt{{2\pi\hbar{{\omega}_{0}}}/{V}}$, $V$ is the field quantization volume ($V$ does not come to final results Lifshitz, Pitaevskii, and Berestetskii (1982)), $\eta$ is the refractive index of the nanoparticle environment Tkalya (2000). We suppose that distances between absorbers and the particle center are much smaller than the external field wavelength, so the phase (taken zero) of the external field is the same for all absorbers and the particle.
We substitute expressions (1), (2) into the sum in Eq (4), drop the fast oscillating terms $\sim e^{\pm 2i\omega_{0}t}$ and represent
$\hat{V}_{int}=\frac{pd}{r^{3}}\sum\limits_{i=1}^{{{N}_{m}}}{\left\{{{{\hat{b}}}^{+}}\left[{{{\hat{\sigma}}}_{i}}-3\left({{{\hat{\sigma}}}_{i}}{{\cos}^{2}}{{\theta}_{i}}+{{{\hat{\sigma}}}_{\bot i}}\cos{{\theta}_{i}}\sin{{\theta}_{i}}\right)\right]+h.c.\right\}}$ (6)
Expression (6) can be simplified at a large number of absorbers $N_{m}\gg 1$. Summing over $i$ we approximate ${{\cos}^{2}}{{\theta}_{i}}\approx 1/2$ and neglect the product ${{{\hat{\sigma}}}_{\bot i}}\cos{{\theta}_{i}}\sin{{\theta}_{i}}$ repeatedly changing the sign with $i$. After such a simplification of Eq. (6) the Hamiltonian (3) written in the interaction picture Landau and Lifshitz (2013) and RWA approximation Sargent, Scully, and Lamb (1974) reads
$H=-{{\hat{a}}^{+}}\left(\hbar{{\Omega}_{p}}\hat{b}+\hbar{{\Omega}_{m}}\sum\limits_{i=1}^{{{N}_{m}}}{{{{\hat{\sigma}}}_{i}}}\right)-\hbar{{\Omega}_{int}}{{\hat{b}}^{+}}\sum\limits_{i=1}^{{{N}_{m}}}{{{{\hat{\sigma}}}_{i}}}+h.c.+\hat{\Gamma}$ (7)
where ${{\Omega}_{p}}={p{{E}_{0}}}/{\hbar}$, ${{\Omega}_{m}}={d{{E}_{0}}}/{\hbar}$ and ${{\Omega}_{int}}={pd}/2\hbar{{r}^{3}}$. The Hamiltonian (7) corresponds to the approximation when the polarization of absorber’s transitions is parallel to the external field polarization.
II.3 Equations of motion
We write Heisenberg equations of motion $i\hbar\dot{\hat{A}}=\left[\hat{A},H\right]$ where $\hat{A}$ is an operator from Hamiltonian (7). The equation for the plasmonic oscillation Bose-operator $\hat{b}$ is
$\dot{\hat{b}}=i{{\Omega}_{p}}\hat{a}+{{\Omega}_{int}}\hat{p}-\Gamma\hat{b}+(\sqrt{2\Gamma}){{\hat{b}}_{in}}$ (8)
where we introduce the polarization operator of all absorbers
$\hat{p}=i\sum\limits_{i=1}^{{{N}_{m}}}{{{{\hat{\sigma}}}_{i}}}.$ (9)
The last two terms are added to Eq. (8) following the input-output theory Courty and Reynaud (1992); Collett and Gardiner (1984), they describe the dissipation of the dipole momentum oscillations with the rate $\Gamma$; ${{\hat{b}}_{in}}$ is the Bose operator of the bath. The stationary solution of Eq. (8) is
$\hat{b}=({i{{\Omega}_{p}}}/{\Gamma})\hat{a}+({{{\Omega}_{int}}}/{\Gamma})\hat{p}+(\sqrt{{2}/{\Gamma}}){{\hat{b}}_{in}}$ (10)
Using well-known properties of Dirac notations Lifshitz, Pitaevskii, and Berestetskii (1982) we come to commutation relations
$\left[{{{\hat{\sigma}}}_{i}},\hat{\sigma}_{i}^{+}\right]={{\hat{N}}_{0i}}-{{\hat{N}}_{1i}}$ (11)
where ${{\hat{N}}_{0i}}={{\left|00\right\rangle}_{i}}{{\left\langle 00\right|}_{i}}$, ${{\hat{N}}_{1i}}={{\left|01\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$ are $i$-th absorber state population operators. Operators with different indexes $i$ commute. Using commutation relations (11), Hamiltonian (7) and polarisation (9) we write the equation of motion
$\dot{\hat{p}}=-\left({{\Omega}_{m}}\hat{a}+{{\Omega}_{int}}\hat{b}\right)\left({{{\hat{N}}}_{0}}-{{{\hat{N}}}_{1}}\right)-\gamma\hat{p}+\hat{F_{p}}$ (12)
where ${{\hat{N}}_{\alpha}}=\sum\limits_{i=1}^{N}{{{{\hat{N}}}_{\alpha i}}}$, $\alpha=0,1$ are operators of populations of states of all absorbers. The last two terms in Eq. (12) describe the absorber polarization decay with the rate $\gamma$ and the noise with the Langevin force $\hat{F_{p}}$ of the absorber polarization baths.
Taking commutation relations $\left[{{{\hat{N}}}_{0i}},{{{\hat{\sigma}}}_{i}}\right]={{\hat{\sigma}}_{i}}$, Hamiltonian (7) and summing over all absorbers we obtain the equation of motion
${{\dot{\hat{N}}}_{0}}=\left({{\Omega}_{m}}{{{\hat{a}}}^{+}}+{{\Omega}_{int}}{{{\hat{b}}}^{+}}\right)\hat{p}+h.c.+{{\hat{N}}_{1}}/\tau+{{\hat{F}}_{0}}$ (13)
where the last two terms are due to the decay of the absorber excited state populations with the decay time $\tau$ and the Langevin force ${{\hat{F}}_{0}}$ corresponding to the population baths. The population operator of the upper states of absorbers is $\hat{N}_{1}=N_{m}-\hat{N}_{0}$.
III The stationary absorption
From here we use the semi-classical approach as, for example, in the laser theory Sargent, Scully, and Lamb (1974). So we neglect fluctuations, drop the bath operators, and replace operators with c-numbers (removing hats from notations). Calculating the semi-classical stationary solution of Eqs. (10) and (12) we find
$p=-\frac{\left({{{\Omega}_{m}}}/{\gamma}+{i{{\Omega}_{int}}{{\Omega}_{p}}}/{\Gamma\gamma}\right)a\left({{N}_{0}}-{{N}_{1}}\right)}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}},$ (14)
where ${{N}_{int}}=\gamma\,\Gamma/\Omega_{int}^{2}$.
Eq. (14) helps us to understand physical processes in the absorber-particle interaction. The stationary polarization of two-level absorbers in the resonant field without the particle is $p=-(\Omega_{m}/\gamma)a(N_{0}-N_{1})$. The particle near absorbers interacts with the external field with the rate $\Omega_{p}$, accepts the polarization, re-emits the field, and, by the re-emitted field, "transfers" the polarization (and the energy) to absorbers with the rate $\Omega_{int}$. Such polarization and the energy transfer is the plasmon-induced resonance energy transfer (PIRET) You, Ramakrishna, and Seideman (2018). PIRET is suppressed by the polarization decay: in the particle with the rate $\Gamma$ and in absorbers with the rate $\gamma$. This is why the PIRET particle-absorber polarization transfer rate is $\sim i\Omega_{int}\Omega_{p}/\Gamma\gamma$, as we see in Eq. (14). The multiplier $i$ in the last expression and in Eq. (14) is due to the well-known $\pi/2$ phase shift at the resonant interaction of oscillators Lifshitz and Landau (1982).
The absorber-particle interaction is a reciprocal process: the particle transfers the polarization to absorbers and vice-versa: absorbers transfer the polarization to the particle through the near field, causing the absorber polarization reduction. Such a reduction is described by the factor $1+\left({{N}_{0}}-{{N}_{1}}\right)/N_{int}$ in the denominator of Eq. (14).
In order to understand the physics and to characterize the stationary absorption we write, using Eq. (13), the energy conservation law: the rate $N_{1}/\tau$ of the energy relaxation from the upper states of absorbers is equal to the sum of the photon absorption rate from the external field $\sim\Omega_{m}$ and the energy exchange rates at the interaction of the particle and absorbers through plasmonic mode $\sim\Omega_{int}$
$\frac{{{N}_{1}}}{\tau}=\left(-{{\Omega}_{m}}+\frac{{{i{\Omega}_{int}\Omega}_{p}}}{\Gamma}\right){{a}^{*}}p+c.c.-\frac{2\Omega_{int}^{2}}{\Gamma}\left\langle{{{\hat{p}}}^{+}}\hat{p}\right\rangle,$ (15)
[…]
$\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}=\hat{N}_{1i}$ (16)
and assuming a large number of absorbers $N_{m}\gg 1$ we write
$\left\langle{{{\hat{p}}}^{+}}\hat{p}\right\rangle={{N}_{1}}+\sum\limits_{i\neq j}{\left\langle\hat{\sigma}_{i}^{+}{{{\hat{\sigma}}}_{j}}\right\rangle}\approx{{N}_{1}}+{{\left|p\right|}^{2}}$ (17)
The term $N_{1}$ in Eq. (17) describes the spontaneous emission, and the second term - a collective emission of absorbers to the plasmonic mode. These terms are similar to the cavity mode spontaneous emission and the collective emission terms in the laser theory André et al. (2019); Protsenko and Uskov (2022). We find from Eqs. (14), (15) and (17)
$\frac{{{N}_{1}}}{\tau}=$ ${\displaystyle\left(\frac{2\Omega_{m}^{2}}{\gamma}+\frac{2\Omega_{p}^{2}}{\Gamma{{N}_{int}}}\right)\frac{n({{N}_{0}}-{{N}_{1}})}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}-}-$
$\frac{2\Omega_{int}^{2}}{\Gamma}\left({{N}_{1}}+{{\left|p\right|}^{2}}\right).$
where $n=|a|^{2}$ is the mean photon number in the volume $V$ of the external field. We express $n/V$ through the external field intensity later.
We introduce the absorption coefficients: ${{g}_{m}}={2\Omega_{m}^{2}}/{\gamma}$ of an absorber, ${{g}_{p}}={2\Omega_{m}^{2}}/{\Gamma}$ of the particle and the spontaneous emission rate of an absorber to the plasmonic mode $\tilde{\gamma}_{sp}=2\Omega_{int}^{2}/\Gamma$. Using Eq. (14) we find
$\frac{2\Omega_{int}^{2}}{\Gamma}{{\left|p\right|}^{2}}=n\frac{\left({{g}_{m}}+{{g}_{p}}/{{N}_{int}}\right){{{\tilde{\gamma}}}_{sp}}}{2\gamma}{{\left(\frac{{{N}_{0}}-{{N}_{1}}}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}\right)}^{2}}$ (19)
and represent Eq. (III) in a compact form
$\frac{{{N}_{1}}}{\tau}=n{{\tilde{g}}_{m}}({{N}_{1}})\frac{{{N}_{0}}-{{N}_{1}}}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}-{{\tilde{\gamma}}_{sp}}{{N}_{1}}$ (20)
where the nonlinear absorption coefficient
${{\tilde{g}}_{m}}({{N}_{1}})=\left({{g}_{m}}+\frac{{{g}_{p}}}{{{N}_{int}}}\right)\left[1-\frac{{{{\tilde{\gamma}}}_{sp}}}{2\gamma}\frac{{{N}_{0}}-{{N}_{1}}}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}\right].$ (21)
With the help of Eqs. (19) – (21) we identify processes affecting the absorber excited states population. The first process is the resonant absorption of the external field with the absorption coefficient $g_{m}$. The second process is the resonant absorption increase due to the nanoparticle near field. An external field energy is absorbed by the particle with the absorption coefficient $g_{p}$ and a part of this energy is transferred to absorbers by the near field with the transfer coefficient $1/N_{int}$. So the linear (i.e. independent on $N_{1}$) part of the absorption coefficient ${{\tilde{g}}_{m}}$ is $g_{m}+g_{p}/N_{int}$. The nonlinear term $\sim\tilde{\gamma}_{sp}$ in Eq. (21) is due to the collective emission of absorbers to plasmonic mode. The reason for the suppression factor $1+(N_{0}-N_{1})/N_{int}$ in the denominator of Eq. (20) is the transfer of the polarization and the energy from absorbers to the particle. The upper states population decays with the rate $\tilde{\gamma}_{sp}$ due to the spontaneous emission to the plasmonic mode, which is described by the last term on the right in Eq. (20).
IV Plasmon-enhanced absorption conditions and bistability
The resonant absorption in plasmonic mode is a nonlinear process. Besides the well-known resonant transition saturation, the nonlinearity is caused by the collective interaction of absorbers and the particle through the plasmonic mode leading to the energy exchange between absorbers and the particle. We will see how such nonlinear processes and the spontaneous emission to the plasmonic mode lead to the bistability in absorption and affect conditions for the best excitation of absorbers. Let us introduce some parameters.
IV.1 Parameters
We express
${{g}_{p}}n=\frac{2\Omega_{p}^{2}n}{\Gamma}=\frac{2n}{\Gamma}{{\left(\frac{p{{E}_{0}}}{\hbar}\right)}^{2}}=\frac{2n}{\Gamma}{{\left(\frac{p}{\hbar}\right)}^{2}}\frac{1}{{{\eta}^{2}}}\frac{2\pi\hbar{{\omega}_{0}}}{V}=\frac{4\pi n\omega_{0}\alpha_{p}}{\eta^{2}V}.$
Here we use ${{{p}^{2}}}/{\hbar\Gamma}={{\alpha}_{p}}$, where ${{\alpha}_{p}}$ is the absolute value of the resonant polarisability of the metal nano-particle Protsenko et al. (2005). The external field intensity is $I=({c\eta}/{4\pi})\left\langle{{{\hat{\vec{E}}}}^{+}}\hat{\vec{E}}\right\rangle={c\hbar{{\omega}_{0}}n}/{V\eta}$. Here and below $c$ is the speed of light in vacuum. We take $I=\hbar{{\omega}_{0}}P$ then the external field intensity is $P={cn}/{V\eta}$ in the number of photons per the unit of area per second. So the number of photons in the external field per the unit of volume is ${n}/{V}={\eta P}/{c}$ and
${{g}_{p}}n={{\alpha}_{p}}\frac{4\pi{{\omega}_{0}}}{\eta}P=\frac{8{{\pi}^{2}}{{\alpha}_{p}}}{\eta{{\lambda}_{0}}}P={{\sigma}_{p}}P,$ (22)
where ${{\lambda}_{0}}=2\pi/{{\omega}_{0}}$ is the absorbing field wavelength in vacuum, $\sigma_{p}$ is the particle resonant absorption cross-section. Similar way we find
${{g}_{m}}n=\frac{8{{\pi}^{2}}{{\alpha}_{m}}}{\eta{{\lambda}_{0}}}P={{\sigma}_{m}}P,$ (23)
where $\alpha_{m}$ is the absolute value of the polarizability and $\sigma_{m}$ is the absorption cross-section of the two-level resonant absorber. It is convenient to represent the resonant absorber’s polarizability
$\alpha_{m}=3\gamma_{sp}/4\eta k_{0}^{3}\gamma,$ (24)
in terms of natural linewidth $\gamma_{sp}$ of the resonant transition Scully and Zubairy (1997); $k_{0}=\omega_{0}/c$ is the absorbed field wave-vector in vacuum. Using the parameters introduced above we express
$\frac{1}{{{N}_{{int}}}}\equiv\frac{\Omega_{{int}}^{2}}{\gamma\Gamma}=\frac{{{p}^{2}}}{\hbar\Gamma}\frac{{{d}^{2}}}{\hbar\gamma}\frac{1}{4{{r}^{6}}}=\frac{{{\alpha}_{p}}{{\alpha}_{m}}}{4{{r}^{6}}},$
$\tilde{\gamma}_{sp}=\frac{2\Omega_{int}^{2}}{\Gamma}=\frac{\alpha_{p}\alpha_{m}}{2r^{6}}\gamma=\frac{2\gamma}{N_{int}}.$
IV.2 Conditions for the absorption enhancement
We derive from the energy conservation law (20) and (21)
${{N}_{1}}={{\sigma}_{eff}}({{N}_{1}})P\tau\left({{N}_{m}}-2{{N}_{1}}\right),$ (25)
where ${{\sigma}_{eff}}({{N}_{1}})$ is a nonlinear (i.e. depended on $N_{1}$) cross-section of absorption
$\sigma_{eff}(N_{1})=\frac{\xi_{\sigma}\sigma_{m}}{[1+(N_{m}-2N_{1})(\alpha_{p}\alpha_{m}/4r^{6})]^{2}}.$ (26)
In Eq. (26) we use $N_{0}+N_{1}=N_{m}$ and introduce the linear absorption cross-section enhancement factor
$\xi_{\sigma}=\frac{1+{\alpha_{p}^{2}}/{4{{r}^{6}}}}{1+\gamma\tau\left({\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}\right)}$ (27)
The nominator of (27) is greater than 1 because of the direct energy transfer from the nanoparticle to absorbing molecules. Such transfer is described by the factor $\alpha_{p}^{2}/4r^{6}$. The denominator of Eq. (27) is greater than 1 because of the spontaneous emission to the LPR mode. Such spontaneous emission reduces the absorption cross-section. $\sigma_{eff}$ is reduced relatively to $\xi_{\sigma}\sigma_{m}$ because of the collective spontaneous emission of absorbers to the LPR mode and the energy transfer back from emitters to the particle. Both these processes make the denominator in Eq. (26) greater than 1.
Conditions, necessary for the absorption enhancement by the LPR follow from Eqs (26), (27). The absorption enhancement is, when $\sigma_{eff}(N_{1})>\sigma_{m}$. Suppose, in the beginning, that the denominator in Eq. (26) is close to 1, so the collective emission to the LPR mode and the energy transfer from absorbers to the nanoparticle are relatively small. This is true at the saturation of the absorbing transitions, when $N_{1}$ is close to $N_{m}/2$, or at a small number $N_{m}$ of absorbers, when
$N_{m}{\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}\leq 1.$ (28)
When Eq. (28) is true, the absorption cross-section is enhanced at $\xi_{\sigma}>1$ which is when
$\alpha_{p}/\alpha_{m}>2\gamma\tau,\hskip 14.22636pt$ (29)
or, if we insert the expression (24) for $\alpha_{m}$ to (29), when $\alpha_{p}>\alpha_{p}^{min}=3\gamma_{sp}\tau/2\eta k_{0}^{3}$. The relation (29) is a necessary condition for the plasmon-stimulated absorption enhancement (found in the semiclassical approach). We note that $2\gamma\tau\geq 1$, so it must be $\alpha_{p}>\alpha_{m}$ for the absorption enhancement. So absorption enhancement is possible when the nanoparticle absorbs the field better than an emitter, which has been mentioned, for example, in Cortés (2017).
However the nanoparticle absorption can not be too large. Otherwise, the absorption by molecules will be suppressed due to their collective emission to the LPR mode and the energy transfer from molecules to the particle. So combining conditions (28) and (29) we obtain the interval of $\alpha_{p}$, where the the absorption enhancement by LPR is possible
$2\gamma\tau\alpha_{m}\equiv\alpha_{p}^{min}\leq\alpha_{p}\leq\alpha_{p}^{max}\equiv 2r^{6}/N_{m}\alpha_{m}.$ (30)
[…]
It is interesting to note that the fact of the absorption enhancement does not depend on the distance $r$ between the absorbers and the particle (as far as absorbers are in the nanoparticle near the field zone). However, the magnitude of the enhancement depends on $r$, as we will see.
IV.3 Bistability in the stationary absorption
Suppose the number $N_{m}$ of absorbing molecules is so large that $N_{m}{\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}>1$. Then the absorption is strongly nonlinear, because of the collective effects, the energy transfer between absorbers and the particle, which makes a significant effect of $N_{1}$-dependent denominator in Eq. (26). The denominator is decreased with $N_{1}$, so the absorption is increased with $N_{1}$ at certain parameter values. As a result the low and the high absorption regimes are possible, they overlap at certain parameters making the bistability in absorption.
In order to investigate the bistability we introduce dimensionless absober-nanoparticle interaction $A={\alpha_{p}\alpha_{m}}{N}_{m}/{4{{r}^{6}}}$, the external field intensity ${{P}_{eff}}=\xi_{\sigma}\sigma_{m}P\tau$ parameters, and consider the upper state population per absorber $n_{1}={N_{1}}/{N_{m}}$. With these parameters we write a cubic equation on $n_{1}$
${{{n}_{1}}}/({1-2{{n}_{1}}})+2{{n}_{1}}A+{{n}_{1}}\left(1-2{{n}_{1}}\right){{A}^{2}}={{P}_{eff}},$ (31)
instead of Eq. (25). Proceeding a standard analysis of the catastrophe theory Poston and Stewart (1996) we find that Eq. (31) has three solutions if the number $N_{m}$ of absorbers is so large that
$A>A_{cr}=8.$ (32)
Note that the condition (32) is opposite to the small collective emission and the small absorber to particle energy transfer condition (28). Thus the collective emission and the absorber to particle energy transfer must be sufficiently large for the bistability.
The bistability domain is restricted by the $P_{\pm}(A)$ curves in $A$, $P_{eff}$ parameter space
$P_{\pm}(A)={{{n}_{\pm}}}/({1-2{{n}_{\pm}}})+2{{n}_{\pm}}A+{{n}_{\pm}}(1-2{{n}_{\pm}}){{A}^{2}}$ (33)
where ${{n}_{\pm}}(A)=\left(3\pm\sqrt{1-8/A}\right)/8$. The bistability domain (33) is shown in Fig. 2.
Figure 2: The absorption bistability is for values of $A$ and $P_{eff}$ between the curves $P_{\pm}(A)$.
V Examples of absorption enhancement and the bistability
In examples we consider a spherical silver nanoparticle with a polymer shell in the water. The refractive index of most polymers is close to the glass refractive index $1.55$ Zhang et al. (2020). The water refractive index is $1.33$. For simplicity, we take the same refractive index $\eta=1.33$ for the nanoparticle shell and the water environment.
Nanoparticle resonant polarizability is calculated in Appendix taking into account the radiative losses Bottcher (1973); Protsenko and Uskov (2012) and the losses due to collisions of the metal electrons with the nanoparticle surface Kik and Brongersma (2007); Protsenko and Uskov (2012). We estimate the LPR wavelength $\lambda_{LPR}=408$ nm and the half-wigth $\Gamma\approx 80$ THz. We take the nanoparticle radius $r_{p}=13$ nm corresponding to the maximum LPR quality factor $Q=29$. Resonant absorbers are on the nanoparticle dielectric shell of the thickness $\Delta r\geq 5$ nm. We take the minimum shell thickness $5$ nm because of a rapid de-population of the absorber excited states (fluorescence quenching) at $\Delta r<5$ nm due to the absorber interaction with the high-order (quadrupole etc.) plasmonic modes Dulkeith et al. (2005). The number of absorbers on the shell is large $N_{m}\geq 10$.
We consider the stationary population $n_{1}$ of the upper level of the absorber resonant transition as a characteristic of the absorption. We compare $n_{1}$ with and without the nano-particle, so we examine the LPR-stimulated absorption enhancement (or suppression) at various numbers of absorbers $N_{m}$, the dielectric shell thickness $\Delta r$, the absorber upper-level de-population time $\tau$ and the absorber transition half-width $\gamma$.
[…]
Figure 3: (a) The stationary upper-level population per absorber versus the normalized external field intensity at various number of absorbers $N_{m}=10$ (curve 1), $80$ (2), and $120$ (3). Curve 4 is for the absorber upper-level population in free space. (b) The ratio $R$ of the emitter upper-level population with and without the nanoparticle (the LPR absorption enhancement) for the same parameters as in Figure 3a. The unstable stationary absorption corresponds to dashed parts of curves 3.
Curves 3 in Figs. 3a, b are for $N_{m}=10$, when $\alpha_{p}/\alpha_{p}^{max}\approx 1.8\sim 1$ and both conditions (28) and (29) (or (30)) are satisfied. We see from curves 1 on Figs. 3 that the LPR absorption is enhanced, relative to the free space absorption, at any input field intensity. The maximum enhancement is about 2.5 –3 times, respectively to free space, at a small field intensity far from the absorber transition saturation.
We increase the number of absorbers. Curves 2 in Figs. 3 are for $N_{m}=80$ so $\alpha_{p}/\alpha_{p}^{max}=14.4\gg 1$ and the condition (28) for a small collective emission and the absorber-to-particle energy transfer is not satisfied. We see the absorption suppression by LPR at small $\sigma_{m}P\tau<1.6$. Otherwise, there is the absorption enhancement for a large $\sigma_{m}P\tau$, when the collective spontaneous emission and the energy transfer are reduced due to the absorber transition saturation. Note that $A=0.97<A_{cr}$ for curves 2, so there is no absorption bistability.
We increase $N_{m}$ more and take $N_{m}=120$ for curves 3 in Figs. 3, so $A=11.6>A_{cr}=8$, the condition (32) is satisfied and we see the bistability in curves 3. The bistability is between regimes with the low and the high absorption. The LPR suppresses the absorption at a small and enhances it at a large external field intensity. The unstable stationary solution corresponds to the dashed parts of curves 3.
Figure 4: $n_{1}$ (a) and $R$ (b) versus the normalized input field intensity for $N_{m}=120$ and different thickness of the nanoparticle shell $\Delta r=5$ nm (curves 1), $6$ nm (2) and 9 nm (3). The rest of the parameters is the same as for Figure 3. The absorber without the nanoparticle corresponds to curve 4. The dashed parts of curves mean the unstable stationary solution. The bistability disappears with the increase of the shell thickness $\Delta r$.
Figure 4 shows $n_{1}$ and $R$ dependencies on the normalized external field intensity $\sigma_{m}P\tau$ for a large number of absorbers $N_{m}=120$ and different thicknesses of the nanoparticle shell $\Delta r=5$ (curves 1), $6$ (2) and $9$ nm (3). We see that the bistability disappears with only 1 nm increase of the nanoparticle shell thickness. For each shell thickness, the absorption is suppressed at small input field intensities and enhanced at high intensities, when the collective emission and the absorber-to-particle energy transfer are suppressed due to the absorber transition saturation.
Let us see what happens with the decrease of the absorber state de-population rate (i.e. the increase $\tau$).
Figure 5: $n_{1}$ (a) and $R$ (b) versus the normalized input field intensity for $N_{m}=10$ and different absorber upper level de-population time $\tau=2$ ps (curves 1), $20$ ps (2) and 50 ps (3). Curve 4 is for an absorber without the nanoparticle. The shell thickness $\Delta r=5$ nm, the rest of parameters is the same as for Figure 3. The upper-level population and the absorption are reduced with the increase of $\tau$.
The absorption enhancement by LPR is reduced with the increase of $\tau$, from curves 1 to curves 3 in Fig 5, because of the growing of the spontaneous emission to the LPR contribution to the population decay.
Suppose we have a large absorber upper state de-population time $\tau=1/2\gamma_{sp}=0.5$ ns; such a de-population is, for example, because of the spontaneous emission to free space with the typical optical domain rate $\gamma_{sp}=1$ GHz Scully and Zubairy (1997). We assume the absorber transition half-width $\gamma=100$ GHz. Then $\alpha_{p}/\alpha_{m}=20<2\gamma\tau=100$ so the condition (29) is not satisfied and there is no absorption enhancement by the LPR at any external field intensity. Meanwhile, the bistability in absorption exists, as shown by curve 3 in Figure 6, when the bistability condition (32) is true.
Figure 6: The upper absorber state population $n_{1}(\sigma_{m}P\tau)$ at the suppression of absorption by the spontaneous emission to LPR mode, when $\tau=0.5$ ns and $\gamma=100$ GHz. The number of absorbers $N_{m}=10$ is for curve 1; $N_{m}=16$ is for curve 2 and $N_{m}=30$ is for curve 3; curve 4 is for the absorber in free space. The bistability condition (32) is satisfied and the bistability in absorption exists for curve 3. The dashed part of curve 3 corresponds to the unstable stationary state.
VI Discussion
An electric field enhancement near plasmonic nanoparticles increases the resonant absorption of atoms, molecules, semiconductor media, etc. The estimation (27) shows the growth of the absorption cross-section near plasmonic particle in a factor $\alpha_{p}^{2}/4r^{6}=9.5$ – without taking into account the spontaneous emission of the absorber to the LPR mode (described by the denominator in Eq. (27)). Spontaneous and collective emissions to the LPR mode significantly reduce the absorption making it, at some conditions, smaller than in the free space. In our examples, the maximum absorption enhancement near plasmonic nanoparticles is about the factor 3 (see curve 1 in Figure 3b), with the spontaneous emission to the LPR mode taken into account.
The maximum absorption enhancement is found for a small external field when the absorbing transition is not saturated. A noticeable absorption enhancement is possible at a fast (on ps time scale) de-population of the upper absorbing state, when the rate of spontaneous emission to the LPR mode is about or less than the de-population rate due to other processes as, for example, the decay of absorber excitation to the non-resonant high-order plasmonic modes Dulkeith et al. (2005). When the population decay is not so fast, the relative suppression of the absorber excitation by the radiation to the LPR is large and we do not have the absorption enhancement, as shown in Figure 6. The condition (29) must be satisfied for the LPR absorption enhancement. Spontaneous and collective emissions to the LPR mode are important factors, they can not be neglected in the resonant absorption near metal nanoparticles.
The spontaneous emission enhancement by the cavity mode is well-known in the laser theory Oraevskii (1994) as a Pursell effect Purcell (1946). The metal nanoparticle is analogous to the cavity for the LPR mode, so such a cavity enhances the spontaneous emission similar to a laser cavity.
When a few resonant absorbers are placed near metal nanoparticles, a collective emission from absorbers to the plasmonic mode is possible. A photon emitted by the atom $i$ is re-absorbed by the atom $j$ and emitted again by the collective emission. This is a Dicke effect Gross and Haroche (1982) with the probability (and, therefore, the rate) larger than for the photon emission by a single atom. The collective emission de-populates the absorber’s upper state. Collective emission to the LPR mode is similar to the collective emission to a lasing mode Leymann et al. (2015); Jahnke et al. (2016); André et al. (2019); Protsenko and Uskov (2022). The collective emission is a nonlinear process, its rate depends on the emitting state populations. In particular, the collective emission is suppressed at the saturation of the emitting transition.
The particle-absorber near-field interaction is a mutual process: the metal nano-particle LPR mode excites the polarization and transfers the energy to the absorber through the near field and vice-versa: absorbers excite the particle polarisation through the near field. Such an energy transfer is the origin of the suppression factor in the denominator in Eq. (14) for the absorber polarisation. The collective emission and the emitter-particle interaction depend on the absorber states population. Such a dependence leads to a positive feedback: the absorption is increased with the upper-level population at certain conditions. The feedback is fast: note the square in the nonlinear denominator in Eq. (26). So, at certain parameter values, the feedback nonlinearity "overtakes" the transition saturation and leads to a decrease of the upper-level population $n_{1}$ with the external field intensity. When $n_{1}$ decreases, the stationary absorption is unstable, see the dashed parts of curves in Figs. 3, 4, and 6. Such anomalous absorption causes absorption bistability. The bistability exists at the condition (32) which can be satisfied at the enhancement (when condition (29) is true) or the suppression (when condition (29) is not true) of absorption by LPR, see Figures 4 and 6. With the LPR absorption enhancement, the bistability is between the absorption suppression and the absorption enhancement regimes, see Figures 3 and 4.
The optical bistability is possible in plasmonic nanolaser with external signal Protsenko et al. (2008). The model of Protsenko et al. (2008) involves a single emitter near the plasmonic nanoparticle and does not include the spontaneous and collective emissions to the LPR mode.
The physical nature of a well-known absorptive optical bistability in a cavity with a two-level absorber (or absorbers) Koch, Schmidt, and Haug (1985) is the resonant transition saturation. This is different from the physics of the bistability in the resonant absorption in the LPR near metal nanoparticle described here. The absorptive bistability Koch, Schmidt, and Haug (1985) is not related to the collective emission and the near-field interaction of absorbers with the metal nanoparticle.
The nanoparticle with resonant emitters is, probably, one of the smallest systems with optical bistability with the size much smaller than the optical wavelength.
Here we carry out the semi-classical analysis of the resonant absorption near a metal nanoparticle. Only the term of the spontaneous emission into the LPR mode in Eq. (17) is introduced according to quantum mechanical relation (16). We note that the term of the spontaneous emission to the lasing mode appears in the laser theory the same way P. R. Rice  and Carmichael (1994); André et al. (2019). More detailed theory of the resonant absorption near plasmonic nanoparticles must include quantum fluctuations of the field and the absorber states populations. Such a quantum analysis is necessary, taking into account a small volume of the plasmonic mode Maier (2006), which keeps only one or a few photons. Such an analysis will be carried out in the future.
VII Conclusion
We investigate the enhancement of the resonant absorption of the electromagnetic field by two-level absorbers (atoms or molecules) near plasmonic nanoparticles, taking into account the spontaneous and collective emission from absorbers to the localized plasmon resonance (LPR) mode of the particle. We derive quantum equations for the particle and the resonant absorber polarisation and the absorbing state populations. We analyze the structure of equations and describe the physical mechanisms of the stationary absorption and the particle-absorber interaction.
We find that the absorption enhancement by the LPR is possible under certain conditions. The absorption suppression is possible as well due to the spontaneous, collective emissions of absorbers to the LPR mode and the energy transfer from absorbers to the particle. We find conditions necessary for the absorption enhancement. One condition is a fast, of the pico-second time scale, de-population of the excited absorber states, for example, by some non-radiative processes.
We find conditions and investigate the bistability in the resonant absorption near the metal nanoparticle. The bistability is caused by the collective emission and the nonlinear interaction between the particle and absorbers through the LPR near field. The bistability appears at a large number of absorbers near the nanoparticle and possible with the absorption enhancement or suppression.
Results can be used everywhere, when the resonant absorption near metal nanoparticles is used, for example, for the plasmon-stimulated photo-catalysis Zhang et al. (2013) or in the plasmonic photo-voltaic Jang et al. (2016).
Acknowledgements.
I.E.P., A.V.U. and N.V.N. acknowledge the Russian Science Foundation (Grant No.20-19-00559) for the support.
Appendix. Nanoparticle resonant polarisability
Polarizability $\alpha_{p}$ of the spherical nanoparticle of the radius $r_{p}$ in the electric field with the wavelength $\lambda$, in the dielectric environment with the refractive index $\eta$ is Protsenko and Uskov (2012)
$\alpha_{p}(\lambda,r_{p})=r_{p}^{3}\frac{\varepsilon_{p}(\lambda,r_{p})/\eta^{2}-1}{[1-iR_{rad}(\lambda,r_{p})][\varepsilon_{p}(\lambda,r_{p})/\eta^{2}+2]}$ (34)
where $\varepsilon_{p}$ is the dielectric function of the material of the nanoparticle and the factor
$R_{rad}(\lambda,r_{p})=\frac{16\pi^{3}}{3}\left(\frac{\eta r_{p}}{\lambda}\right)^{3}\frac{\varepsilon_{p}(\lambda,r_{p})/\eta^{2}-1}{\varepsilon_{p}(\lambda,r_{p})/\eta^{2}+2}$ (35)
describe the radiation losses.
The nanoparticle material dielectric function $\varepsilon_{p}(\lambda,r_{p})$ takes into account the losses due to collisions of the nanoparticle electrons with the nanoparticle boundary Protsenko and Uskov (2012); Uskov et al. (2014)
$\varepsilon_{p}(\lambda,r_{p})=\varepsilon_{0}-\frac{(\lambda/\lambda_{p})^{2}}{(1+i\lambda/\lambda_{c})(r_{c}/r_{p}+1)}.$ (36)
[…]

### v2

@@
-
+Igor E. Protsenko Alexander V. Uskov Nikolay V. Nikonorov
 Abstract
-We find that the spontaneous and collective emissions have a large influence on the excitation of two-level absorbers (atoms, molecules) interacting in resonance with the plasmonic mode near the metal nanoparticle.
-The spontaneous and collective emissions restrict the absorption enhancement by plasmonic mode and make the enhancement possible only with a fast, on a pico-second time scale, population relaxation of the upper absorbing states.
-Conditions for the maximum of the plasmon-enhanced absorption in the presence of spontaneous and collective emissions are found.
-We show that the nanoparticle-absorber interaction and the collective emission cause the bistability in the plasmon-enhanced absorption.
-I Introduction
-The localized plasmon resonances (LPR) excitation leads to the electric field enhancement near metal nanoparticles Maier (2007); Pelton, Aizpurua, and Bryant (2008); Amendola et al.
-(2017); Tanabe (2008).
-Such a "plasmonic" field enhancement finds various applications in optical sensors Li, Cushing, and Wu (2015), bio-medicine Liao, Nehl, and Hafner (2006), photovoltaics Jang et al.
-(2016), photo-catalysis Zhang et al.
-(2013) and other areas of physics and chemistry Maier (2007); Barbillon (2019).
-A new branch of chemistry - plasmonic chemistry was developed concerning the chemical reactions Zhan et al.
-(2023); Cortés (2017) stimulated by the electric field enhanced near plasmonic nanoparticles.
+We find that the spontaneous and collective emissions have a strong influence on the excitation of two-level absorbers (atoms, molecules) interacting in resonance with the plasmonic mode near the metal nanoparticle.
+The spontaneous and collective emissions limit the absorption enhancement by the plasmonic mode and make the enhancement possible only with a fast, picosecond population relaxation of the upper absorbing states.
+Conditions for the maximum of plasmon-enhanced absorption in the presence of spontaneous and collective emissions are found.
+The nonlinearity in the nanoparticle-absorber interaction and in collective emission causes the bistability in the plasmon-enhanced absorption at high external field intensities and the plasmonic mode excitation.
+keywords:
+plasmonics , nanoparticle , absorption
+PACS:
+78.67.Bf , 73.20.Mf , 78.67.Qa , 73.22.Dj
+journal: Photonics and nanostructures …
+[…]
+[inst1]organization=P.N.Lebedev Physical Institute of the RAS,addressline=Leninsky prospect 53, city=Moscow, postcode=119991, country=Russia
+[…]
+[inst2]organization=ITMO University,addressline=Kronverksky Pr. 49, city=St. Petersburg, postcode=197101, country=Russia
+1 Introduction
+Excitation of localised plasmon resonances (LPR) leads to electric field enhancement in the vicinity of metal nanoparticles [1, 2, 3, 4].
+Such a ”plasmonic” field enhancement finds various applications in optical sensors [5], bio-medicine [6], photovoltaics [7], photocatalysis [8] and other areas of physics and chemistry [1, 9].
+A new branch of chemistry - plasmonic chemistry - has been developed around the chemical reactions [10, 11] stimulated by the electric field enhancement near plasmonic nanoparticles.
 A photochemical reaction is initiated by the absorption of light and the creation of transient molecules in excited states whose properties differ greatly from the original molecules.
-These new chemical species can fall apart, change to new structures, combine other molecules, or transfer electrons, hydrogen atoms, protons, or their electronic excitation energy to other molecules Fleming, Krueger, and Longworth (2023).
-So the optimization of the plasmonic photo-chemical reactions is important with the searching for conditions of the maximum field and absorption enhancement near metal nanoparticles Kelly et al.
-(2003).
-This paper contributes to such an optimization.
-A large electric field enhancement is well-known near the sharp features of nanostructures, as "hot spots" Díaz-Núñez et al.
-(2019), or in a small gap between nanoparticles Kim et al.
-(2010).
-However, it may be difficult to make a sufficiently large number of "hot spots" or small gaps in nanostructures.
-So in this paper, we do not consider the field enhancement by "hot spots" or in gaps between nanoparticles.
-We restrict ourselves by the absorption enhancement in the near field of metal nanoparticles with the excited LPR.
-We will pay attention on the spontaneous and collective emissions, which significantly affect the resonant absorption near plasmonic nanoparticles and cause new physical effects.
-Spontaneous emission near metal nanoparticles has been investigated, for example, in Blanco and García de Abajo (2004).
-Collective effects as a super-radiance have been studied for the resonant absorbers, as atoms or molecules, in the near zone of the plasmonic particle Shahbazyan and Pustovit (2011); Protsenko and Uskov (2015); Protsenko et al.
-(2017).
-However, the absorption enhancement near the nanoparticle with an account for the spontaneous and collective emissions deserves more studies.
-The influence of spontaneous emission and collective effects on absorption is important for plasmonic chemistry, particularly, in the relation with recently studied plasmon-induced direct resonance energy transfer (PIRET) processes You, Ramakrishna, and Seideman (2018).
-A part of studies of the plasmon-enhanced absorption with the spontaneous and collective emissions to the LPR mode, with the direct energy transfer from the plasminic nanoparticle to the absorbing molecule, is done in the present paper.
-Here we use the model for the resonant absorption near plasmonic nanoparticles similar to the laser model.
-The plasmonic nanoparticle makes a cavity for the electric field similar to the laser cavity, as in the plasmonic nanolaser Protsenko et al.
-(2005, 2008); Protsenko (2012).
-A part of our approach is similar to the analysis of the spontaneous emission and collective effects in lasers André et al.
-(2019); Protsenko and Uskov (2022).
+These new chemical species can fall apart, change into new structures, combine other molecules, or transfer electrons, hydrogen atoms, protons, or their electronic excitation energy to other molecules [12].
+Therefore, it is important to optimise the plasmonic photochemical reactions by searching for conditions of maximum field and absorption enhancement near metal nanoparticles [13].
+This work contributes to such optimization.
+A large electric field enhancement is known to occur near the sharp features of nanostructures, known as ’hot spots’ [14], or in a small gap between nanoparticles [15].
+A plasmonic gap field mode has been shown to have a high potential for achieving ultra-strong light absorption, the magnitude of which can be relatively easily controlled [16, 17, 18, 19].
+In this paper we consider, as a first approximation, a much simpler system, limited to the near-field absorption enhancement of the isolated metal nanoparticle with the excited LPR.
+The approach and results of the present simple model will help to analyse the perspective but more complicated problem of absorption enhancement in plasmonic gaps.
+We will focus on the spontaneous and collective emissions, which significantly affect the resonant absorption near plasmonic nanoparticles and lead to new physical effects.
+Spontaneous emission near metal nanoparticles has been studied in many papers, for example in [20].
+Collective effects in the form of superradiance have been studied for the resonant absorbers, such as atoms or molecules, in the vicinity of the plasmonic particle [21, 22, 23].
+However, the absorption enhancement in the vicinity of the nanoparticle, taking into account the spontaneous and collective emissions, deserves further study.
+The influence of spontaneous emission and collective effects on absorption is important for plasmonic chemistry, especially in the context of the recently studied plasmon-induced direct resonance energy transfer (PIRET) processes [24].
+Part of the study of plasmon-enhanced absorption with spontaneous and collective emission in the LPR mode, with direct energy transfer from the plasmonic nanoparticle to the absorbing molecule, is carried out in the present work.
+Here we use the model for resonant absorption near plasmonic nanoparticles, similar to the laser model.
+The plasmonic nanoparticle forms a cavity for the electric field similar to the laser cavity as in the plasmonic nanolaser [25, 26, 27].
+Part of our approach is similar to the analysis of spontaneous emission and collective effects in lasers [28, 29].
 The plasmonic mode provides good conditions for collective effects in the resonant absorption.
-Collective effects in the resonant medium are significant in the low-quality cavity André et al.
-(2019); Protsenko and Uskov (2022); Piccione, Felicetti, and Bellomo (2022).
-The LPR mode has a low quality, appropriate for collective effects.
-The particle surface is a good place for the resonant absorbers (atoms, molecules, q-dots, etc.).
-In this paper, we show that the spontaneous emission into the LPR mode and collective interaction of absorbers through the plasmonic electric field significantly affect the absorber excitation in the vicinity of metal nanoparticles and lead to bistability in the excitation.
-We will find conditions for the best absorber excitation and bistability.
-These results are important for all branches of plasmonic chemistry, photovoltaic, and all applications related to the resonant excitation of absorbers near plasmonic nanoparticles.
-We derive basic equations quantum-mechanically as Heisenberg operator equations.
-In the present paper, we solve these equations in the semi-classical approximation as it is common, for example, in the laser theory Sargent, Scully, and Lamb (1974).
-Fully quantum analysis of the equations will be done in the future.
-We carry out calculations analytically.
-The analytical approach lets to find expressions illustrating the physical processes at the absorber excitation and the influence of the plasmonic particle on important parameters such as the polarization and the population of states of absorbers.
-We consider the increase of the absorber excited state population as a criterion of absorption enhancement by the LPR.
-In the next section, we describe the model, Hamiltonian, and equations of motion.
-The expressions, characterizing the stationary absorption, as the absorption coefficients and the energy balance equation, are presented in section III.
-Several physical processes at the resonant absorption near the metal nanoparticle are explained with the help of equations found in section III.
-Section IV shows conditions for the best plasmon-enhanced absorption and describes the bistability.
-Examples of the absorption enhancement and the bistability are presented in section V at typical values of the nanoparticle and the absorber parameters.
-Results are discussed in section VI, and the paper is finalized in the conclusion section VII.
-II The model, Hamiltonian and equations of motion
-II.1 The model
-Suppose, a metal nanoparticle interacts with $i=1...N_{m}$ identical absorbers, as molecules or atoms, being on the distance $r$ from the particle center on the dielectric shell as shown in Fig.
+Collective effects in the resonant medium are significant in the low quality cavity [28, 29, 30].
+The LPR mode is of low quality.
+The particle surface is a good location for the resonant absorbers (atoms, molecules, q-dots, etc.).
+In this paper, we show that the spontaneous emission in the LPR mode and the collective interaction of absorbers through the plasmonic electric field significantly affect the absorber excitation in the vicinity of metal nanoparticles and lead to bistability in the excitation.
+We will find conditions for absorber excitation enhancement and bistability.
+These results are important for all branches of plasmonic chemistry, photovoltaics and all applications related to resonant excitation of absorbers near plasmonic nanoparticles.
+We derive the fundamental equations quantum mechanically as Heisenberg operator equations.
+In the present work, we solve these equations in the semi-classical approximation, as it is used for example in laser theory [31].
+A full quantum analysis of the equations will be performed in the future.
+We perform analytical calculations.
+The analytical approach allows us to find expressions that illustrate the physical processes at the excitation of the absorber and the influence of the plasmonic particle on important parameters such as the polarisation and the population of states of the absorber.
+We consider the increase in the absorber excited state population as a criterion for absorption enhancement by the LPR.
+We describe the interaction of the dipole absorbers (molecules) with the dipole mode of the spherical plasmonic nanoparticle.
+According to Mie theory [32], the external field of the nanoparticle dipole mode is the field of the dipole in the nanoparticle center.
+Thus, the nanoparticle-absorber interaction is the dipole-dipole interaction as long as the high-order plasmonic mode excitation is neglected.
+We assume the resonance between the absorber transition and the frequency of the plasmonic dipole mode.
+The resonant dipole-dipole interaction at a short absorber-nanoparticle distance causes the well-known Förster nonradiative resonance energy transfer (FRET) [33].
+Thus, in our work, the effect of collective and single absorber spontaneous emission in the plasmonic mode on the FRET process is highlighted.
+We assume that the molecules are located on the nanoparticle dielectric shell with a thickness of at least 5 nm, which is sufficient to avoid strong absorber fluorescence quenching [34].
+We assume that the nanoparticle radius is $13$ nm, so the total distance of $18$ nm between the nanoparticle center and the absorbers on the shell is much smaller than the optical wavelength of the resonant field $\sim 1~{}\mu$m, see the scheme in Fig.
 1.
-The absorber transition frequency is resonant with the dipole LPR mode of the nanoparticle.
+We therefore use the quasi-static approximation for the absorber-nanoparticle dipole-dipole interaction.
+We take into account the nonradiative and radiative [35] losses of the nanoparticle, including the losses associated with the collisions of the metal electrons with the nanoparticle surface [35, 36].
+Thus, such a model based on the dipole-dipole interaction of the absorbers with the nanoparticle is a good approximation for our purposes.
+We note that more sophisticated models, such as macroscopic QED for quantum nanophotonics [37] and other approaches as discussed in the review [38], can be used for advanced investigation of the spontaneous and collective emission in plasmonic nanostructures outlined in this paper.
+A relatively simple model given below is useful for finding similarities in the collective emission phenomena in small lasers, LEDs and resonant absorber plasmonic structures.
+In the next section, we describe the model, the Hamiltonian, and the equations of motion.
+The expressions characterising the steady-state absorption, such as the absorption coefficients and the energy balance equation, are presented in section 3.
+Several physical processes in the resonant absorption near the metal nanoparticle are explained with the help of the equations found in section 3.
+Section 4 shows the conditions for the best plasmon-enhanced absorption and describes the bistability.
+Examples of the absorption enhancement and the bistability are shown in section 5 for typical values of the nanoparticle and the absorber parameters.
+The results are discussed in section 6 and the paper is concluded in section 7.
+2 The model, Hamiltonian and equations of motion
+2.1 The model
+Suppose, a metal nanoparticle interacts with $i=1...N_{m}$ identical absorbers, as molecules or atoms, located at the distance $r$ from the particle center on the dielectric shell as shown in Fig.
+1.
+The transition frequency of the absorber is in resonance with the dipole LPR mode of the nanoparticle.
 Figure 1: Coordinate system, the nanoparticle 1 and the resonant absorber $i$ (a molecule or an atom) in the external electric field $\hat{\vec{E}}$.
-$r$ is the distance, and $\vec{e}_{ri}$ is the unit vector from the nanoparticle center to the absorber.
-$\left|00\right>_{i}$ is the absorber ground state.
-The dipole momentum with the matrix element $d$ for the absorber transition from the excited state $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to $\left|00\right>_{i}$ is parallel to the unit vector $\vec{e}$ of the external field polarization (or to the vector $\vec{e}_{i}$).
+$r$ is the distance, and $\vec{e}_{ri}$ is the unit vector from the centre of the nanoparticle to the absorber.
+$\left|00\right>_{i}$ is the ground state of the absorber.
+The dipole momentum with the matrix element $d$ for the absorber transition from the excited state $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to $\left|00\right>_{i}$ is parallel to the unit vector $\vec{e}$ of the external field polarisation (or to the vector $\vec{e}_{i}$).
 The absorber state with the transition dipole momentum parallel to the axis $OO^{\prime}$ is not excited; such transition does not interact with the external and the nanoparticle fields.
-The insert shows the resonant absorbers randomly distributed on the nanoparticle shell.
-Such resonance can be achieved for the nanoparticle of appropriate shape Kelly et al. (2003). For simplicity, we consider a spherical nanoparticle.
-The particle and absorbers interact with the external monochromatic electric field of the amplitude $\hat{\vec{E}}$ linearly polarized along the unit vector $\vec{e}$.
-The field $\hat{\vec{E}}$ is resonant with the LPR and the absorber transitions.
-We consider the $i$-th absorber in the Cartesian coordinate system with the origin in the particle center.
-One coordinate axis is parallel to $\vec{e}$, another axis with the unit vector $\vec{e}_{i}$ is in the plane made by $\vec{e}$ and the unit vector $\vec{e}_{ri}$ directed from the nanoparticle center to the i-th absorber.
-The angle $\theta_{i}$ is between $\vec{e}$ and $\vec{e}_{i}$, and the third coordinate axis is along the line $OO^{\prime}$.
-Two excited ${{\left\langle 10\right|}_{i}}$, ${{\left\langle 01\right|}_{i}}$ and the ground ${{\left\langle 00\right|}_{i}}$ states of i-th absorber participate in the interaction with electric fields near the nanoparticle.
-Such absorber states are, for example, the lowest unoccupied molecular orbital (LUMO) and the highest occupied molecular orbital (HOMO) Fukui, Yonezawa, and Shingu (2004).
-We will calculate the population of excited states of absorbers and examine how the interaction between the particle and absorbers affects the population.
-We suppose a large number $N_{m}\gg 1$ of absorbers randomly distributed on the nanoparticle shell as shown on the insert in Fig.
+The inset shows the resonant absorbers randomly distributed on the nanoparticle shell.
+Such resonance can be achieved for the nanoparticle of appropriate shape [13].
+For simplicity, we will consider a spherical nanoparticle.
+A metallic nanoparticle with resonant absorbers/emitters on the dielectric shell is a well-known object in plasmonic theory and experiments.
+It has been considered and realised, for example, in the plasmonic nano-lasers [39, 40, 27], for the studies of the plasmonic superradiance [41, 22] and, most recently, in the discovery of the role of electron spillout in strong coupling of the plasmonic nanoparticle with the excitonic shell [42].
+The particle and the absorbers interact with the external monochromatic electric field of the amplitude $\hat{\vec{E}}$ linearly polarised along the unit vector $\vec{e}$.
+The field $\hat{\vec{E}}$ is in resonance with the LPR and the absorber transitions.
+We consider the $i$-th absorber in the Cartesian coordinate system with the origin at the particle centre.
+One coordinate axis is parallel to $\vec{e}$, another axis with the unit vector $\vec{e}_{i}$ is in the plane formed by $\vec{e}$ and the unit vector $\vec{e}_{ri}$ points from the nanoparticle centre to the i-th absorber.
+The angle $\theta_{i}$ is between $\vec{e}$ and $\vec{e}_{i}$, and the third axis of coordinate is along the line $OO^{\prime}$.
+Two excited ${{\left\langle 10\right|}_{i}}$, ${{\left\langle 01\right|}_{i}}$ and the ground ${{\left\langle 00\right|}_{i}}$ states of the i-th absorber participate interact with electric fields near the nanoparticle.
+Such absorber states are, for example, the lowest unoccupied molecular orbital (LUMO) and the highest occupied molecular orbital (HOMO) [43].
+We will calculate the population of excited states of the absorbers and study how the interaction between the particle and the absorbers affects the population.
+We assume a large number $N_{m}\gg 1$ of absorbers randomly distributed on the nanoparticle shell, as shown in the inset in Fig.
 1.
-In the case of $N_{m}\gg 1$, the absorber-absorber interaction through the near field leads to only the self-broadening of the absorber’s transitions Royer (1975).
-We include the self-broadening in the absorber transition width $\gamma$ and do not describe the absorber-absorber interaction explicitly.
-II.2 Hamiltonian
-Following Protsenko et al. (2005) we describe the nanoparticle as a dipole making harmonic oscillations in the resonant external electric field. We introduce the nanoparticle dipole momentum operator
+In the case of $N_{m}\gg 1$, the absorber-absorber interaction through the near-field leads only to the self-broadening of the absorber’s transitions [44].
+We include the self-broadening in the absorber transition width $\gamma$ and do not explicitly describe the absorber-absorber interaction.
+2.2 Hamiltonian
+Following [25], we describe the nanoparticle as a dipole that undergoes harmonic oscillations in the resonant external electric field. We introduce the nanoparticle dipole momentum operator
 $\hat{\vec{p}}=p\vec{e}\left(\hat{b}{{e}^{-i{{\omega}_{0}}t}}+{{{\hat{b}}}^{+}}{{e}^{i{{\omega}_{0}}t}}\right)$ (1)
 where $p$ is the dipole momentum matrix element, $\hat{b}$ is a Bose-operator describing collective (plasmonic) oscillations of the nanoparticle electrons in the electric field.
-Later we express $p$ through the nanoparticle polarizability Protsenko et al.
-(2005).
-For symmetry reasons, we suppose the same directions for the nanoparticle dipole and the external field polarization.
-The dipole momentum operator of the $i$-th absorber resonant transitions is
-${{\hat{\vec{d}}}_{i}}=d\left[\left({{{\hat{\sigma}}}_{i}}\vec{e}+{{{\hat{\sigma}}}_{\bot i}}{{{\vec{e}}}_{i}}\right){{e}^{-i{{\omega}_{0}}t}}+h.c.\right].$ (2)
-Here $d$ is the dipole transition matrix element; absorber operators in Dirac notations ${{\hat{\sigma}}_{i}}={{\left|00\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$, (or ${{{\hat{\sigma}}}_{\bot i}}={{\left|00\right\rangle}_{i}}{{\left\langle 10\right|}_{i}}$) are transition operators from the high $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to the low $\left|00\right>_{i}$ states of the absorber $i$ parallel to $\vec{e}$, (or $\vec{e}_{i}$): ${{{\hat{\sigma}}}_{i}}\left|01\right>_{i}=\left|00\right>_{i}$ and ${{{\hat{\sigma}}}_{\perp i}}\left|10\right>_{i}=\left|00\right>_{i}$.
-Hamiltonian of the particle and $N_{m}$ absorbers in the electric field $\hat{\vec{E}}$ is
+Later we express $p$ in terms of the polarisability of the nanoparticles [25].
+For reasons of symmetry, we assume the same directions for the nanoparticle dipole and the external field polarisation.
+The dipole momentum operator of the $i$th absorber resonance transition is
+${{\hat{\vec{d}}}_{i}}=d\left[\left({{{\hat{\sigma}}}_{i}}\vec{e}+{{{\hat{\sigma}}}_{\bot i}}{{{\vec{e}}}_{i}}\right){{e}^{-i{{\omega}_{0}}t}}+h.c.\right],$ (2)
+where $d$ is the dipole transition matrix element; absorber operators in Dirac notation ${{\hat{\sigma}}_{i}}={{\left|00\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$, (or ${{{\hat{\sigma}}}_{\bot i}}={{\left|00\right\rangle}_{i}}{{\left\langle 10\right|}_{i}}$) are transition operators from the high $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to the low $\left|00\right>_{i}$ states of the absorber $i$ parallel to $\vec{e}$, (or $\vec{e}_{i}$): ${{{\hat{\sigma}}}_{i}}\left|01\right>_{i}=\left|00\right>_{i}$ and ${{{\hat{\sigma}}}_{\perp i}}\left|10\right>_{i}=\left|00\right>_{i}$.
+The Hamiltonian of the particle and $N_{m}$ absorbers in the electric field $\hat{\vec{E}}$ is
 $H={{H}_{0}}-\left(\hat{\vec{p}}+\sum\limits_{i=1}^{N}{{{{\hat{\vec{d}}}}_{i}}}\right)\hat{\vec{E}}+{{\hat{V}}_{int}}+\hat{\Gamma}$ (3)
-The first term in (3) describes free oscillations of dipoles; the second term is for the dipole interaction of the particle and absorbers with the external field; the third term ${{\hat{V}}_{int}}$ describes the interaction of the particle and absorbers through the nanoparticle near field (the LPR field) and the last term in (3) describes the dissipation and the interaction of the particle and absorbers with baths.
-We suppose that the distance $r$ between absorbers and the particle center is much smaller than the LPR wavelength.
-So the particle-absorber interaction ${{\hat{V}}_{int}}$ is a quasi-static dipole-dipole interaction.
-We take the dipole-dipole quasi-static interaction energy from Landau and Lifshitz (1982), replace classical dipoles with operators, and write
+The first term in (3) describes the free oscillations of the dipoles; the second term is for the dipole interaction of the particle and absorbers with the external field; the third term ${{\hat{V}}_{int}}$ describes the interaction of the particle and absorbers through the nanoparticle near field (the LPR field), and the last term in (3) describes the dissipation and the interaction of the particle and absorbers with baths.
+We assume that the distance $r$ between the absorbers and the particle centre is much smaller than the LPR wavelength.
+Thus the particle-absorber interaction ${{\hat{V}}_{int}}$ is a quasi-static dipole-dipole interaction.
+We take the quasi-static dipole-dipole interaction energy from [45], replace the classical dipoles by operators, and write
 ${{\hat{V}}_{int}}=\frac{1}{{{r}^{3}}}\sum\limits_{i=1}^{{{N}_{m}}}{\left[\hat{\vec{p}}{{{\hat{\vec{d}}}}_{i}}-3(\hat{\vec{p}}{{{\vec{e}}}_{ri}})({{{\hat{\vec{d}}}}_{i}}{{{\vec{e}}}_{ri}})\right]},$ (4)
-where ${{{\vec{e}}}_{ri}}$ is the unit vector along the direction from the nanoparticle center to i-th absorber as shown in Fig. 1.
+where ${{{\vec{e}}}_{ri}}$ is the unit vector along the direction from the nanoparticle centre to the ith absorber, as shown in Fig. 1.
 The external field is
 $\hat{\vec{E}}={{E}_{0}}\vec{e}\left(\hat{a}{{e}^{-i{{\omega}_{0}t}}}+{{{\hat{a}}}^{+}}{{e}^{i{{\omega}_{0}}t}}\right)$ (5)
-where $\hat{a}$ is the field amplitude Bose-operator, ${{E}_{0}}=\eta^{-1}\sqrt{{2\pi\hbar{{\omega}_{0}}}/{V}}$, $V$ is the field quantization volume ($V$ does not come to final results Lifshitz, Pitaevskii, and Berestetskii (1982)), $\eta$ is the refractive index of the nanoparticle environment Tkalya (2000).
-We suppose that distances between absorbers and the particle center are much smaller than the external field wavelength, so the phase (taken zero) of the external field is the same for all absorbers and the particle.
-We substitute expressions (1), (2) into the sum in Eq (4), drop the fast oscillating terms $\sim e^{\pm 2i\omega_{0}t}$ and represent
+where $\hat{a}$ is the field amplitude Bose operator, ${{E}_{0}}=\eta^{-1}\sqrt{{2\pi\hbar{{\omega}_{0}}}/{V}}$, $V$ is the field quantization volume ($V$ is not used in the final results [46]), $\eta$ is the refractive index of the nanoparticle environment [47].
+We assume that the distances between the absorbers and the particle centre are much smaller than the wavelength of the external field, so that the phase (assumed to be zero) of the external field is the same for all absorbers and the particle.
+Substituting the expressions (1), (2) into the sum in Eq (4), dropping the fast oscillating terms $\sim e^{\pm 2i\omega_{0}t}$, we get
 $\hat{V}_{int}=\frac{pd}{r^{3}}\sum\limits_{i=1}^{{{N}_{m}}}{\left\{{{{\hat{b}}}^{+}}\left[{{{\hat{\sigma}}}_{i}}-3\left({{{\hat{\sigma}}}_{i}}{{\cos}^{2}}{{\theta}_{i}}+{{{\hat{\sigma}}}_{\bot i}}\cos{{\theta}_{i}}\sin{{\theta}_{i}}\right)\right]+h.c.\right\}}$ (6)
-Expression (6) can be simplified at a large number of absorbers $N_{m}\gg 1$.
-Summing over $i$ we approximate ${{\cos}^{2}}{{\theta}_{i}}\approx 1/2$ and neglect the product ${{{\hat{\sigma}}}_{\bot i}}\cos{{\theta}_{i}}\sin{{\theta}_{i}}$ repeatedly changing the sign with $i$.
+Expression (6) can be simplified for a large number of absorbers $N_{m}\gg 1$.
+Summing over $i$, we approximate ${{\cos}^{2}}{{\theta}_{i}}\approx 1/2$ and neglect the product ${{{\hat{\sigma}}}_{\bot i}}\cos{{\theta}_{i}}\sin{{\theta}_{i}}$ repeatedly changing the sign with $i$.
 After such a simplification of Eq.
-(6) the Hamiltonian (3) written in the interaction picture Landau and Lifshitz (2013) and RWA approximation Sargent, Scully, and Lamb (1974) reads
+(6), the Hamiltonian (3) written in the interaction picture [48] and RWA approximation [31] becomes
 $H=-{{\hat{a}}^{+}}\left(\hbar{{\Omega}_{p}}\hat{b}+\hbar{{\Omega}_{m}}\sum\limits_{i=1}^{{{N}_{m}}}{{{{\hat{\sigma}}}_{i}}}\right)-\hbar{{\Omega}_{int}}{{\hat{b}}^{+}}\sum\limits_{i=1}^{{{N}_{m}}}{{{{\hat{\sigma}}}_{i}}}+h.c.+\hat{\Gamma}$ (7)
 where ${{\Omega}_{p}}={p{{E}_{0}}}/{\hbar}$, ${{\Omega}_{m}}={d{{E}_{0}}}/{\hbar}$ and ${{\Omega}_{int}}={pd}/2\hbar{{r}^{3}}$.
-The Hamiltonian (7) corresponds to the approximation when the polarization of absorber’s transitions is parallel to the external field polarization.
-II.3 Equations of motion
-We write Heisenberg equations of motion $i\hbar\dot{\hat{A}}=\left[\hat{A},H\right]$ where $\hat{A}$ is an operator from Hamiltonian (7).
-The equation for the plasmonic oscillation Bose-operator $\hat{b}$ is
+The Hamiltonian (7) corresponds to the approximation when the polarisation of the absorber transitions is parallel to the polarisation of the external field.
+2.3 Equations of motion
+We write Heisenberg the equations of motion $i\hbar\dot{\hat{A}}=\left[\hat{A},H\right]$ where $\hat{A}$ is an operator from the Hamiltonian (7).
+The equation for the plasmonic oscillation Bose operator $\hat{b}$ is
 $\dot{\hat{b}}=i{{\Omega}_{p}}\hat{a}+{{\Omega}_{int}}\hat{p}-\Gamma\hat{b}+(\sqrt{2\Gamma}){{\hat{b}}_{in}}$ (8)
-where we introduce the polarization operator of all absorbers
+where we introduce the polarisation operator of all absorbers
 $\hat{p}=i\sum\limits_{i=1}^{{{N}_{m}}}{{{{\hat{\sigma}}}_{i}}}.$ (9)
 The last two terms are added to Eq.
-(8) following the input-output theory Courty and Reynaud (1992); Collett and Gardiner (1984), they describe the dissipation of the dipole momentum oscillations with the rate $\Gamma$; ${{\hat{b}}_{in}}$ is the Bose operator of the bath.
+(8) according to the input-output theory [49, 50], they describe the dissipation of the dipole momentum oscillations with the rate $\Gamma$; ${{\hat{b}}_{in}}$ is the Bose operator of the bath.
 The stationary solution of Eq.
@@
 $\hat{b}=({i{{\Omega}_{p}}}/{\Gamma})\hat{a}+({{{\Omega}_{int}}}/{\Gamma})\hat{p}+(\sqrt{{2}/{\Gamma}}){{\hat{b}}_{in}}$ (10)
-Using well-known properties of Dirac notations Lifshitz, Pitaevskii, and Berestetskii (1982) we come to commutation relations
+Using well-known properties of Dirac notations [46], we arrive at the commutation relations
 $\left[{{{\hat{\sigma}}}_{i}},\hat{\sigma}_{i}^{+}\right]={{\hat{N}}_{0i}}-{{\hat{N}}_{1i}}$ (11)
-where ${{\hat{N}}_{0i}}={{\left|00\right\rangle}_{i}}{{\left\langle 00\right|}_{i}}$, ${{\hat{N}}_{1i}}={{\left|01\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$ are $i$-th absorber state population operators.
-Operators with different indexes $i$ commute.
-Using commutation relations (11), Hamiltonian (7) and polarisation (9) we write the equation of motion
+where ${{\hat{N}}_{0i}}={{\left|00\right\rangle}_{i}}{{\left\langle 00\right|}_{i}}$, ${{\hat{N}}_{1i}}={{\left|01\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$ are ith absorber state population operators.
+Operators with different indices $i$ are commute.
+Using the commutation relations (11), Hamiltonian (7), and polarisation (9) we write the equation of motion
 $\dot{\hat{p}}=-\left({{\Omega}_{m}}\hat{a}+{{\Omega}_{int}}\hat{b}\right)\left({{{\hat{N}}}_{0}}-{{{\hat{N}}}_{1}}\right)-\gamma\hat{p}+\hat{F_{p}}$ (12)
@@
 The last two terms in Eq.
-(12) describe the absorber polarization decay with the rate $\gamma$ and the noise with the Langevin force $\hat{F_{p}}$ of the absorber polarization baths.
-Taking commutation relations $\left[{{{\hat{N}}}_{0i}},{{{\hat{\sigma}}}_{i}}\right]={{\hat{\sigma}}_{i}}$, Hamiltonian (7) and summing over all absorbers we obtain the equation of motion
+(12) describe the absorber polarisation decay with the rate $\gamma$ and the noise with the Langevin force $\hat{F_{p}}$ of the absorber polarisation baths.
+Taking the commutation relations $\left[{{{\hat{N}}}_{0i}},{{{\hat{\sigma}}}_{i}}\right]={{\hat{\sigma}}_{i}}$, Hamiltonian (7), and summing over all absorbers we obtain the equation of motion
 ${{\dot{\hat{N}}}_{0}}=\left({{\Omega}_{m}}{{{\hat{a}}}^{+}}+{{\Omega}_{int}}{{{\hat{b}}}^{+}}\right)\hat{p}+h.c.+{{\hat{N}}_{1}}/\tau+{{\hat{F}}_{0}}$ (13)
-where the last two terms are due to the decay of the absorber excited state populations with the decay time $\tau$ and the Langevin force ${{\hat{F}}_{0}}$ corresponding to the population baths.
-The population operator of the upper states of absorbers is $\hat{N}_{1}=N_{m}-\hat{N}_{0}$.
-III The stationary absorption
-From here we use the semi-classical approach as, for example, in the laser theory Sargent, Scully, and Lamb (1974).
-So we neglect fluctuations, drop the bath operators, and replace operators with c-numbers (removing hats from notations).
+where the last two terms are due to the decay of the absorber excited state populations with the decay time $\tau$ and the Langevin force ${{\hat{F}}_{0}}$ corresponding to the population bath.
+The population operator of the absorber upper states is $\hat{N}_{1}=N_{m}-\hat{N}_{0}$.
+3 The stationary absorption
+From here we use the semi-classical approach as, for example, in the laser theory [31].
+We neglect fluctuations, drop the bath operators, and replace the operators with c-numbers (removing the hats from the notations).
 Calculating the semi-classical stationary solution of Eqs.
@@
 Eq.
-(14) helps us to understand physical processes in the absorber-particle interaction.
-The stationary polarization of two-level absorbers in the resonant field without the particle is $p=-(\Omega_{m}/\gamma)a(N_{0}-N_{1})$.
-The particle near absorbers interacts with the external field with the rate $\Omega_{p}$, accepts the polarization, re-emits the field, and, by the re-emitted field, "transfers" the polarization (and the energy) to absorbers with the rate $\Omega_{int}$.
-Such polarization and the energy transfer is the plasmon-induced resonance energy transfer (PIRET) You, Ramakrishna, and Seideman (2018).
-PIRET is suppressed by the polarization decay: in the particle with the rate $\Gamma$ and in absorbers with the rate $\gamma$.
-This is why the PIRET particle-absorber polarization transfer rate is $\sim i\Omega_{int}\Omega_{p}/\Gamma\gamma$, as we see in Eq.
+(14) helps us to understand the physical processes in the absorber-particle interaction.
+The stationary polarisation of two-level absorbers in the resonant field without the particle is $p=-(\Omega_{m}/\gamma)a(N_{0}-N_{1})$.
+The particle near the absorbers interacts with the external field at the rate $\Omega_{p}$, accepts the polarisation, re-emits the field, and through the re-emitted field ”transfers” the polarisation (and the energy) to absorbers with the rate $\Omega_{int}$.
+Such polarisation and the energy transfer is called plasmon induced resonance energy transfer (PIRET) [24].
+PIRET is suppressed by the polarisation decay: in the particle at the rate $\Gamma$ and in the absorbers at the rate $\gamma$.
+Therefore, the PIRET particle-absorber polarisation transfer rate is $\sim i\Omega_{int}\Omega_{p}/\Gamma\gamma$, as we see in Eq.
 (14).
 The multiplier $i$ in the last expression and in Eq.
-(14) is due to the well-known $\pi/2$ phase shift at the resonant interaction of oscillators Lifshitz and Landau (1982).
-The absorber-particle interaction is a reciprocal process: the particle transfers the polarization to absorbers and vice-versa: absorbers transfer the polarization to the particle through the near field, causing the absorber polarization reduction.
+(14) is due to the well-known $\pi/2$ phase shift in the resonant interaction of oscillators [51].
+The interaction between the absorber and the particle is a two-way process: the particle transfers the polarisation to the absorber and vice versa: the absorber transfers the polarisation to the particle through the near field, causing a reduction in the polarisation of the absorber.
 Such a reduction is described by the factor $1+\left({{N}_{0}}-{{N}_{1}}\right)/N_{int}$ in the denominator of Eq.
 (14).
-In order to understand the physics and to characterize the stationary absorption we write, using Eq.
-(13), the energy conservation law: the rate $N_{1}/\tau$ of the energy relaxation from the upper states of absorbers is equal to the sum of the photon absorption rate from the external field $\sim\Omega_{m}$ and the energy exchange rates at the interaction of the particle and absorbers through plasmonic mode $\sim\Omega_{int}$
+To understand the physics and to characterise the stationary absorption we write the energy conservation law using Eq.
+(13): the rate $N_{1}/\tau$ of the energy relaxation from the upper states of the absorbers is equal to the sum of the photon absorption rate from the external field $\sim\Omega_{m}$ and the energy exchange rates at the interaction of the particle with the absorbers through the plasmonic mode $\sim\Omega_{int}$
 $\frac{{{N}_{1}}}{\tau}=\left(-{{\Omega}_{m}}+\frac{{{i{\Omega}_{int}\Omega}_{p}}}{\Gamma}\right){{a}^{*}}p+c.c.-\frac{2\Omega_{int}^{2}}{\Gamma}\left\langle{{{\hat{p}}}^{+}}\hat{p}\right\rangle,$ (15)
@@
 The term $N_{1}$ in Eq.
-(17) describes the spontaneous emission, and the second term - a collective emission of absorbers to the plasmonic mode.
-These terms are similar to the cavity mode spontaneous emission and the collective emission terms in the laser theory André et al.
-(2019); Protsenko and Uskov (2022).
-We find from Eqs.
-(14), (15) and (17)
-$\frac{{{N}_{1}}}{\tau}=$ ${\displaystyle\left(\frac{2\Omega_{m}^{2}}{\gamma}+\frac{2\Omega_{p}^{2}}{\Gamma{{N}_{int}}}\right)\frac{n({{N}_{0}}-{{N}_{1}})}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}-}-$
-$\frac{2\Omega_{int}^{2}}{\Gamma}\left({{N}_{1}}+{{\left|p\right|}^{2}}\right).$
-where $n=|a|^{2}$ is the mean photon number in the volume $V$ of the external field. We express $n/V$ through the external field intensity later.
+(17) describes the spontaneous emission, and the second term describes a collective emission of absorbers to the plasmonic mode.
+These terms are similar to the cavity mode spontaneous emission and the collective emission terms in the laser theory [28, 29].
+From Eqs.
+(14), (15) and (17) we find
+$\frac{{{N}_{1}}}{\tau}=\displaystyle\left(\frac{2\Omega_{m}^{2}}{\gamma}+\frac{2\Omega_{p}^{2}}{\Gamma{{N}_{int}}}\right)\frac{n({{N}_{0}}-{{N}_{1}})}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}-\frac{2\Omega_{int}^{2}}{\Gamma}\left({{N}_{1}}+{{\left|p\right|}^{2}}\right)$ (18)
+where $n=|a|^{2}$ is the mean number of photons in the volume $V$ of the external field. We will later express $n/V$ in terms of the external field intensity.
 We introduce the absorption coefficients: ${{g}_{m}}={2\Omega_{m}^{2}}/{\gamma}$ of an absorber, ${{g}_{p}}={2\Omega_{m}^{2}}/{\Gamma}$ of the particle and the spontaneous emission rate of an absorber to the plasmonic mode $\tilde{\gamma}_{sp}=2\Omega_{int}^{2}/\Gamma$.
@@
 $\frac{2\Omega_{int}^{2}}{\Gamma}{{\left|p\right|}^{2}}=n\frac{\left({{g}_{m}}+{{g}_{p}}/{{N}_{int}}\right){{{\tilde{\gamma}}}_{sp}}}{2\gamma}{{\left(\frac{{{N}_{0}}-{{N}_{1}}}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}\right)}^{2}}$ (19)
-and represent Eq. (III) in a compact form
+and represent Eq. (18) in a compact form
 $\frac{{{N}_{1}}}{\tau}=n{{\tilde{g}}_{m}}({{N}_{1}})\frac{{{N}_{0}}-{{N}_{1}}}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}-{{\tilde{\gamma}}_{sp}}{{N}_{1}}$ (20)
@@
 ${{\tilde{g}}_{m}}({{N}_{1}})=\left({{g}_{m}}+\frac{{{g}_{p}}}{{{N}_{int}}}\right)\left[1-\frac{{{{\tilde{\gamma}}}_{sp}}}{2\gamma}\frac{{{N}_{0}}-{{N}_{1}}}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}\right].$ (21)
-With the help of Eqs.
-(19) – (21) we identify processes affecting the absorber excited states population.
+Using Eqs.
+(19) – (21) we identify processes that affect the absorber excited state population.
 The first process is the resonant absorption of the external field with the absorption coefficient $g_{m}$.
-The second process is the resonant absorption increase due to the nanoparticle near field.
-An external field energy is absorbed by the particle with the absorption coefficient $g_{p}$ and a part of this energy is transferred to absorbers by the near field with the transfer coefficient $1/N_{int}$.
-So the linear (i.e.
-independent on $N_{1}$) part of the absorption coefficient ${{\tilde{g}}_{m}}$ is $g_{m}+g_{p}/N_{int}$.
-The nonlinear term $\sim\tilde{\gamma}_{sp}$ in Eq.
+The second process is the resonant absorption enhancement due to the nanoparticle near field.
+An external field energy is absorbed by the particle with the absorption coefficient $g_{p}$ and a part of this energy is transferred to the absorbers by the near field with the transfer coefficient $1/N_{int}$.
+Thus the linear (i.e.
+independent of $N_{1}$) part of the absorption coefficient ${{\tilde{g}}_{m}}$ is $g_{m}+g_{p}/N_{int}$.
+The non-linear term $\sim\tilde{\gamma}_{sp}$ in Eq.
 (21) is due to the collective emission of absorbers to plasmonic mode.
 The reason for the suppression factor $1+(N_{0}-N_{1})/N_{int}$ in the denominator of Eq.
-(20) is the transfer of the polarization and the energy from absorbers to the particle.
-The upper states population decays with the rate $\tilde{\gamma}_{sp}$ due to the spontaneous emission to the plasmonic mode, which is described by the last term on the right in Eq.
+(20) is due to the transfer of the polarisation and the energy from the absorbers to the particle.
+The upper state population decays at the rate $\tilde{\gamma}_{sp}$ due to the spontaneous emission to the plasmonic mode, which is described by the last term on the right in Eq.
 (20).
-IV Plasmon-enhanced absorption conditions and bistability
-The resonant absorption in plasmonic mode is a nonlinear process.
-Besides the well-known resonant transition saturation, the nonlinearity is caused by the collective interaction of absorbers and the particle through the plasmonic mode leading to the energy exchange between absorbers and the particle.
-We will see how such nonlinear processes and the spontaneous emission to the plasmonic mode lead to the bistability in absorption and affect conditions for the best excitation of absorbers.
+4 Plasmon-enhanced absorption conditions and bistability
+Resonant absorption in the plasmonic mode is a non-linear process.
+In addition to the well-known resonant transition saturation, the nonlinearity is caused by by the collective interaction of the absorbers and the particle through the plasmonic mode, leading to an energy exchange between the absorbers and the particle.
+We will see how such nonlinear processes and the spontaneous emission to the plasmonic mode lead to the bistability of the absorption and affect the conditions for the best excitation of the absorbers.
 Let us introduce some parameters.
-IV.1 Parameters
+4.1 Parameters
 We express
 ${{g}_{p}}n=\frac{2\Omega_{p}^{2}n}{\Gamma}=\frac{2n}{\Gamma}{{\left(\frac{p{{E}_{0}}}{\hbar}\right)}^{2}}=\frac{2n}{\Gamma}{{\left(\frac{p}{\hbar}\right)}^{2}}\frac{1}{{{\eta}^{2}}}\frac{2\pi\hbar{{\omega}_{0}}}{V}=\frac{4\pi n\omega_{0}\alpha_{p}}{\eta^{2}V}.$
-Here we use ${{{p}^{2}}}/{\hbar\Gamma}={{\alpha}_{p}}$, where ${{\alpha}_{p}}$ is the absolute value of the resonant polarisability of the metal nano-particle Protsenko et al.
-(2005).
+Here we use ${{{p}^{2}}}/{\hbar\Gamma}={{\alpha}_{p}}$, where ${{\alpha}_{p}}$ is the absolute value of the resonant polarisability of the metal nanoparticle [25].
 The external field intensity is $I=({c\eta}/{4\pi})\left\langle{{{\hat{\vec{E}}}}^{+}}\hat{\vec{E}}\right\rangle={c\hbar{{\omega}_{0}}n}/{V\eta}$.
 Here and below $c$ is the speed of light in vacuum.
-We take $I=\hbar{{\omega}_{0}}P$ then the external field intensity is $P={cn}/{V\eta}$ in the number of photons per the unit of area per second.
+If we take $I=\hbar{{\omega}_{0}}P$ then the external field intensity is $P={cn}/{V\eta}$ in the number of photons per the unit of area per second.
 So the number of photons in the external field per the unit of volume is ${n}/{V}={\eta P}/{c}$ and
 ${{g}_{p}}n={{\alpha}_{p}}\frac{4\pi{{\omega}_{0}}}{\eta}P=\frac{8{{\pi}^{2}}{{\alpha}_{p}}}{\eta{{\lambda}_{0}}}P={{\sigma}_{p}}P,$ (22)
-where ${{\lambda}_{0}}=2\pi/{{\omega}_{0}}$ is the absorbing field wavelength in vacuum, $\sigma_{p}$ is the particle resonant absorption cross-section. Similar way we find
+where ${{\lambda}_{0}}=2\pi/{{\omega}_{0}}$ is the wavelength of the absorbing field in vacuum, $\sigma_{p}$ is the absorption cross section of the particle resonance. Similarly, we find
 ${{g}_{m}}n=\frac{8{{\pi}^{2}}{{\alpha}_{m}}}{\eta{{\lambda}_{0}}}P={{\sigma}_{m}}P,$ (23)
-where $\alpha_{m}$ is the absolute value of the polarizability and $\sigma_{m}$ is the absorption cross-section of the two-level resonant absorber.
-It is convenient to represent the resonant absorber’s polarizability
+where $\alpha_{m}$ is the absolute value of the polarisability and $\sigma_{m}$ is the absorption cross section of the two-level resonant absorber.
+It is convenient to express the polarisability of the resonant absorber
 $\alpha_{m}=3\gamma_{sp}/4\eta k_{0}^{3}\gamma,$ (24)
-in terms of natural linewidth $\gamma_{sp}$ of the resonant transition Scully and Zubairy (1997); $k_{0}=\omega_{0}/c$ is the absorbed field wave-vector in vacuum.
-Using the parameters introduced above we express
+in terms of the natural linewidth $\gamma_{sp}$ of the resonant transition [52]; $k_{0}=\omega_{0}/c$ is the absorbed field wave vector in vacuum. Using the parameters introduced above, we express
 $\frac{1}{{{N}_{{int}}}}\equiv\frac{\Omega_{{int}}^{2}}{\gamma\Gamma}=\frac{{{p}^{2}}}{\hbar\Gamma}\frac{{{d}^{2}}}{\hbar\gamma}\frac{1}{4{{r}^{6}}}=\frac{{{\alpha}_{p}}{{\alpha}_{m}}}{4{{r}^{6}}},$
 $\tilde{\gamma}_{sp}=\frac{2\Omega_{int}^{2}}{\Gamma}=\frac{\alpha_{p}\alpha_{m}}{2r^{6}}\gamma=\frac{2\gamma}{N_{int}}.$
-IV.2 Conditions for the absorption enhancement
-We derive from the energy conservation law (20) and (21)
+4.2 Conditions for the absorption enhancement
+From the energy conservation law (20) and (21) we derive
 ${{N}_{1}}={{\sigma}_{eff}}({{N}_{1}})P\tau\left({{N}_{m}}-2{{N}_{1}}\right),$ (25)
-where ${{\sigma}_{eff}}({{N}_{1}})$ is a nonlinear (i.e. depended on $N_{1}$) cross-section of absorption
+where ${{\sigma}_{eff}}({{N}_{1}})$ is a non-linear (i.e. dependent on $N_{1}$) absorption cross section
 $\sigma_{eff}(N_{1})=\frac{\xi_{\sigma}\sigma_{m}}{[1+(N_{m}-2N_{1})(\alpha_{p}\alpha_{m}/4r^{6})]^{2}}.$ (26)
-In Eq. (26) we use $N_{0}+N_{1}=N_{m}$ and introduce the linear absorption cross-section enhancement factor
-$\xi_{\sigma}=\frac{1+{\alpha_{p}^{2}}/{4{{r}^{6}}}}{1+\gamma\tau\left({\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}\right)}$ (27)
-The nominator of (27) is greater than 1 because of the direct energy transfer from the nanoparticle to absorbing molecules.
+In Eq. (26) we use $N_{0}+N_{1}=N_{m}$ and introduce the linear absorption cross section enhancement factor
+$\xi_{\sigma}=\frac{1+{\alpha_{p}^{2}}/{4{{r}^{6}}}}{1+\gamma\tau\left({\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}\right)}.$ (27)
+The nominator of (27) is greater than 1 because of the direct energy transfer from the nanoparticle to the absorbing molecules.
 Such transfer is described by the factor $\alpha_{p}^{2}/4r^{6}$.
@@
 (27) is greater than 1 because of the spontaneous emission to the LPR mode.
-Such spontaneous emission reduces the absorption cross-section.
-$\sigma_{eff}$ is reduced relatively to $\xi_{\sigma}\sigma_{m}$ because of the collective spontaneous emission of absorbers to the LPR mode and the energy transfer back from emitters to the particle.
-Both these processes make the denominator in Eq.
+Such spontaneous emission reduces the absorption cross section.
+$\sigma_{eff}$ is reduced relative to $\xi_{\sigma}\sigma_{m}$ due to the collective spontaneous emission of absorbers to the LPR mode and the energy transfer back from absorbers to the particle.
+These two processes make the denominator in Eq.
 (26) greater than 1.
-Conditions, necessary for the absorption enhancement by the LPR follow from Eqs (26), (27).
-The absorption enhancement is, when $\sigma_{eff}(N_{1})>\sigma_{m}$.
-Suppose, in the beginning, that the denominator in Eq.
-(26) is close to 1, so the collective emission to the LPR mode and the energy transfer from absorbers to the nanoparticle are relatively small.
-This is true at the saturation of the absorbing transitions, when $N_{1}$ is close to $N_{m}/2$, or at a small number $N_{m}$ of absorbers, when
+The conditions, necessary for the absorption enhancement by the LPR follow from Eqs (26), (27).
+The absorption enhancement is present when $\sigma_{eff}(N_{1})>\sigma_{m}$.
+Assume at the beginning, that the denominator in Eq.
+(26) is close to 1, so that the collective emission to the LPR mode and the energy transfer from the absorbers to the nanoparticle are relatively small.
+This is the case at the saturation of the absorbing transitions, when $N_{1}$ is close to $N_{m}/2$, or at a small number of $N_{m}$ of absorbers, when
 $N_{m}{\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}\leq 1.$ (28)
-When Eq. (28) is true, the absorption cross-section is enhanced at $\xi_{\sigma}>1$ which is when
+If Eq. (28) holds, the absorption cross section is enhanced at $\xi_{\sigma}>1$ which is when
 $\alpha_{p}/\alpha_{m}>2\gamma\tau,\hskip 14.22636pt$ (29)
-or, if we insert the expression (24) for $\alpha_{m}$ to (29), when $\alpha_{p}>\alpha_{p}^{min}=3\gamma_{sp}\tau/2\eta k_{0}^{3}$.
+or, if we insert the expression (24) for $\alpha_{m}$ into (29), when $\alpha_{p}>\alpha_{p}^{min}=3\gamma_{sp}\tau/2\eta k_{0}^{3}$.
 The relation (29) is a necessary condition for the plasmon-stimulated absorption enhancement (found in the semiclassical approach).
-We note that $2\gamma\tau\geq 1$, so it must be $\alpha_{p}>\alpha_{m}$ for the absorption enhancement.
-So absorption enhancement is possible when the nanoparticle absorbs the field better than an emitter, which has been mentioned, for example, in Cortés (2017).
-However the nanoparticle absorption can not be too large.
-Otherwise, the absorption by molecules will be suppressed due to their collective emission to the LPR mode and the energy transfer from molecules to the particle.
-So combining conditions (28) and (29) we obtain the interval of $\alpha_{p}$, where the the absorption enhancement by LPR is possible
+Note that $2\gamma\tau\geq 1$, so $\alpha_{p}>\alpha_{m}$ is required for absorption enhancement.
+So absorption enhancement is possible if the nanoparticle absorbs the field better than an absorber, as mentioned for example in [11].
+However, the absorption of the nanoparticles must not be too high.
+Otherwise, the absorption by the molecules will be suppressed due to their collective emission to the LPR mode and the energy transfer from the molecules to the particle.
+Thus, by combining conditions (28) and (29) we obtain the interval of $\alpha_{p}$, where the absorption enhancement by LPR is possible
 $2\gamma\tau\alpha_{m}\equiv\alpha_{p}^{min}\leq\alpha_{p}\leq\alpha_{p}^{max}\equiv 2r^{6}/N_{m}\alpha_{m}.$ (30)
 […]
-It is interesting to note that the fact of the absorption enhancement does not depend on the distance $r$ between the absorbers and the particle (as far as absorbers are in the nanoparticle near the field zone).
-However, the magnitude of the enhancement depends on $r$, as we will see.
-IV.3 Bistability in the stationary absorption
-Suppose the number $N_{m}$ of absorbing molecules is so large that $N_{m}{\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}>1$.
-Then the absorption is strongly nonlinear, because of the collective effects, the energy transfer between absorbers and the particle, which makes a significant effect of $N_{1}$-dependent denominator in Eq.
+It is interesting to note that the fact of the absorption enhancement does not depend on the distance $r$ between the absorbers and the particle (as long as the absorbers are in the nanoparticle near the field zone).
+However, the magnitude of the enhancement does depend on $r$, as we will see.
+4.3 Bistability in the stationary absorption
+Suppose the number $N_{m}$ of absorbing molecules is such that $N_{m}{\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}>1$.
+Then the absorption is strongly non-linear due to the collective effects, the energy transfer between the absorbers and the particle, which has a significant effect of the $N_{1}$-dependent denominator in Eq.
 (26).
-The denominator is decreased with $N_{1}$, so the absorption is increased with $N_{1}$ at certain parameter values.
-As a result the low and the high absorption regimes are possible, they overlap at certain parameters making the bistability in absorption.
-In order to investigate the bistability we introduce dimensionless absober-nanoparticle interaction $A={\alpha_{p}\alpha_{m}}{N}_{m}/{4{{r}^{6}}}$, the external field intensity ${{P}_{eff}}=\xi_{\sigma}\sigma_{m}P\tau$ parameters, and consider the upper state population per absorber $n_{1}={N_{1}}/{N_{m}}$.
-With these parameters we write a cubic equation on $n_{1}$
+The denominator decreases with $N_{1}$, so the absorption increases with $N_{1}$ for certain parameter values.
+As a result, the low and the high absorption regimes are possible, overlapping at certain parameters, creating the bistability in the absorption.
+To study the bistability we introduce the dimensionless nanoparticle-absorber interaction $A={\alpha_{p}\alpha_{m}}{N}_{m}/{4{{r}^{6}}}$, the external field intensity ${{P}_{eff}}=\xi_{\sigma}\sigma_{m}P\tau$ parameters, and consider the upper state population per absorber $n_{1}={N_{1}}/{N_{m}}$.
+With these parameters we write a cubic equation for $n_{1}$
 ${{{n}_{1}}}/({1-2{{n}_{1}}})+2{{n}_{1}}A+{{n}_{1}}\left(1-2{{n}_{1}}\right){{A}^{2}}={{P}_{eff}},$ (31)
-instead of Eq. (25). Proceeding a standard analysis of the catastrophe theory Poston and Stewart (1996) we find that Eq. (31) has three solutions if the number $N_{m}$ of absorbers is so large that
+instead of Eq. (25). Following a standard analysis of the catastrophe theory [53], we find that Eq. (31) has three solutions when the number $N_{m}$ of absorbers is large enough that
 $A>A_{cr}=8.$ (32)
-Note that the condition (32) is opposite to the small collective emission and the small absorber to particle energy transfer condition (28).
+Note that the condition (32) is opposite to the small collective emission and the small absorber-to-particle energy transfer condition (28).
 Thus the collective emission and the absorber to particle energy transfer must be sufficiently large for the bistability.
-The bistability domain is restricted by the $P_{\pm}(A)$ curves in $A$, $P_{eff}$ parameter space
+The bistability domain is bounded by the $P_{\pm}(A)$ curves in $A$, $P_{eff}$ parameter space
 $P_{\pm}(A)={{{n}_{\pm}}}/({1-2{{n}_{\pm}}})+2{{n}_{\pm}}A+{{n}_{\pm}}(1-2{{n}_{\pm}}){{A}^{2}}$ (33)
 where ${{n}_{\pm}}(A)=\left(3\pm\sqrt{1-8/A}\right)/8$. The bistability domain (33) is shown in Fig. 2.
-Figure 2: The absorption bistability is for values of $A$ and $P_{eff}$ between the curves $P_{\pm}(A)$.
-V Examples of absorption enhancement and the bistability
-In examples we consider a spherical silver nanoparticle with a polymer shell in the water.
-The refractive index of most polymers is close to the glass refractive index $1.55$ Zhang et al.
-(2020).
-The water refractive index is $1.33$.
-For simplicity, we take the same refractive index $\eta=1.33$ for the nanoparticle shell and the water environment.
-Nanoparticle resonant polarizability is calculated in Appendix taking into account the radiative losses Bottcher (1973); Protsenko and Uskov (2012) and the losses due to collisions of the metal electrons with the nanoparticle surface Kik and Brongersma (2007); Protsenko and Uskov (2012).
-We estimate the LPR wavelength $\lambda_{LPR}=408$ nm and the half-wigth $\Gamma\approx 80$ THz.
-We take the nanoparticle radius $r_{p}=13$ nm corresponding to the maximum LPR quality factor $Q=29$.
-Resonant absorbers are on the nanoparticle dielectric shell of the thickness $\Delta r\geq 5$ nm.
-We take the minimum shell thickness $5$ nm because of a rapid de-population of the absorber excited states (fluorescence quenching) at $\Delta r<5$ nm due to the absorber interaction with the high-order (quadrupole etc.) plasmonic modes Dulkeith et al.
-(2005).
+Figure 2: Absorption bistability for values of $A$ and $P_{eff}$ lie between the curves $P_{\pm}(A)$.
+5 Examples of absorption enhancement and the bistability
+As an example, consider a spherical silver nanoparticle with a polymer shell in the water.
+The refractive index of most polymers is close to the refractive index of glass $1.55$ [54].
+The refractive index of water is $1.33$.
+For simplicity, we assume the same refractive index $\eta=1.33$ for the nanoparticle shell and the water environment.
+The resonant polarisability of the nanoparticles is calculated in the Appendix, taking into account the radiative losses [55, 35] and the losses due to collisions of the metal electrons with the nanoparticle surface [56, 35].
+We estimate the LPR wavelength $\lambda_{LPR}=408$ nm and the half width $\Gamma\approx 80$ THz.
+We take the nanoparticle radius $r_{p}=13$ nm, which corresponds to the maximum LPR quality factor $Q=29$.
+The resonant absorbers are located on the nanoparticle dielectric shell of the thickness $\Delta r\geq 5$ nm.
+We chose a minimum shell thickness, because of a rapid increase in absorber fluorescence quenching near metal nanoparticles has been observed at smaller $\Delta r$ [34, 57].
+Strong quenching at small $\Delta r$ can be caused by several mechanisms, besides the dipole-dipole interaction, such as a nanometal surface energy transfer (NSET) [58].
+Taking $\Delta r\geq 5$ nm we can neglect by the nanoparticle-absorber’s interactions, but the dipole-dipole one, to keep our system as simple as possible.
 The number of absorbers on the shell is large $N_{m}\geq 10$.
-We consider the stationary population $n_{1}$ of the upper level of the absorber resonant transition as a characteristic of the absorption.
-We compare $n_{1}$ with and without the nano-particle, so we examine the LPR-stimulated absorption enhancement (or suppression) at various numbers of absorbers $N_{m}$, the dielectric shell thickness $\Delta r$, the absorber upper-level de-population time $\tau$ and the absorber transition half-width $\gamma$.
-[…]
-Figure 3: (a) The stationary upper-level population per absorber versus the normalized external field intensity at various number of absorbers $N_{m}=10$ (curve 1), $80$ (2), and $120$ (3).
+We consider the stationary population $n_{1}$ of the upper level of the absorber resonance transition as a property of the absorption.
+We compare $n_{1}$ with and without the nanoparticle, so we examine the LPR-stimulated absorption enhancement (or suppression) at different numbers of absorbers $N_{m}$, the dielectric shell thickness $\Delta r$, the absorber upper level depopulation time $\tau$ and the absorber transition half width $\gamma$.
+Suppose we have a fast depopulation of the upper absorber states with a population decay time $\tau=2$ ps.
+Such a depopulation determines the absorber transition half-width so $\gamma=1/2\tau$.
+We take the parameter values written above and find that the condition (29) for the LPR absorption enhancement is satisfied: $\alpha_{p}/\alpha_{m}=100>2\gamma\tau=1$.
+Figure 3a shows $n_{1}$, Figure 3b - the ratio $R=n_{1}/n_{1}^{(0)}$ at different $N_{m}$, as a function of the normalised external field intensity $\sigma_{m}P\tau$; $n_{1}^{(0)}$ is an absorber upper level population calculated without the nanoparticle.
+At current parameter values, $\sigma_{m}P\tau=1$ corresponds to an external field intensity 0.54 MWt/cm2.
+Such intensities can be achieved, for example, in a pulsed laser regime with pulse duration $\sim 1$ ns $\gg\tau=2$ ps, so that we have a quasi-stationary population of absorber states.
+The external field energy is about 0.5 mJ/pulse at a pulse cross section of 1 cm2.
+Figure 3: (a) The steady-state upper-level population per absorber versus the normalised external field intensity for different numbers of absorbers $N_{m}=10$ (curve 1), $80$ (2), and $120$ (3).
 Curve 4 is for the absorber upper-level population in free space.
-(b) The ratio $R$ of the emitter upper-level population with and without the nanoparticle (the LPR absorption enhancement) for the same parameters as in Figure 3a.
-The unstable stationary absorption corresponds to dashed parts of curves 3.
+(b) The ratio $R$ of the absorber upper-level population with and without the nanoparticle (the LPR absorption enhancement) for the same parameters as in Figure 3a.
+The unstable steady-state absorption corresponds to the dashed parts of curves 3.
 Curves 3 in Figs.
 3a, b are for $N_{m}=10$, when $\alpha_{p}/\alpha_{p}^{max}\approx 1.8\sim 1$ and both conditions (28) and (29) (or (30)) are satisfied.
-We see from curves 1 on Figs.
+We can see from curves 1 in Figs.
 3 that the LPR absorption is enhanced, relative to the free space absorption, at any input field intensity.
-The maximum enhancement is about 2.5 –3 times, respectively to free space, at a small field intensity far from the absorber transition saturation.
+The maximum enhancement is about 2.5 –3 times that of free space at a small field intensity far from the absorber transition saturation.
 We increase the number of absorbers.
 Curves 2 in Figs.
-3 are for $N_{m}=80$ so $\alpha_{p}/\alpha_{p}^{max}=14.4\gg 1$ and the condition (28) for a small collective emission and the absorber-to-particle energy transfer is not satisfied.
+3 are for $N_{m}=80$ so $\alpha_{p}/\alpha_{p}^{max}=14.4\gg 1$ and the condition (28) for a small collective emission and energy transfer from absorber to particle is not satisfied.
 We see the absorption suppression by LPR at small $\sigma_{m}P\tau<1.6$.
-Otherwise, there is the absorption enhancement for a large $\sigma_{m}P\tau$, when the collective spontaneous emission and the energy transfer are reduced due to the absorber transition saturation.
+Otherwise, there is the absorption enhancement for a large $\sigma_{m}P\tau$, when the collective spontaneous emission and the energy transfer are reduced due to the saturation of the absorber transition.
 Note that $A=0.97<A_{cr}$ for curves 2, so there is no absorption bistability.
-We increase $N_{m}$ more and take $N_{m}=120$ for curves 3 in Figs.
+We increase $N_{m}$ further and take $N_{m}=120$ for curves 3 in Figs.
 3, so $A=11.6>A_{cr}=8$, the condition (32) is satisfied and we see the bistability in curves 3.
-The bistability is between regimes with the low and the high absorption.
-The LPR suppresses the absorption at a small and enhances it at a large external field intensity.
-The unstable stationary solution corresponds to the dashed parts of curves 3.
-Figure 4: $n_{1}$ (a) and $R$ (b) versus the normalized input field intensity for $N_{m}=120$ and different thickness of the nanoparticle shell $\Delta r=5$ nm (curves 1), $6$ nm (2) and 9 nm (3).
-The rest of the parameters is the same as for Figure 3.
+The bistability is between regimes with the low and the high absorption regimes.
+The LPR suppresses the absorption at low external field intensities and enhances it at high external field intensities.
+The unstable steady state solution corresponds to the dashed parts of curves 3.
+Figure 4: $n_{1}$ (a) and $R$ (b) versus the normalised input field intensity for $N_{m}=120$ and different nanoparticle shell thickness $\Delta r=5$ nm (curves 1), $6$ nm (2) and 9 nm (3).
+The rest of the parameters are the same as in Figure 3.
 The absorber without the nanoparticle corresponds to curve 4.
-The dashed parts of curves mean the unstable stationary solution.
+The dashed parts of the curves represent the unstable steady-state solution.
 The bistability disappears with the increase of the shell thickness $\Delta r$.
-Figure 4 shows $n_{1}$ and $R$ dependencies on the normalized external field intensity $\sigma_{m}P\tau$ for a large number of absorbers $N_{m}=120$ and different thicknesses of the nanoparticle shell $\Delta r=5$ (curves 1), $6$ (2) and $9$ nm (3).
-We see that the bistability disappears with only 1 nm increase of the nanoparticle shell thickness.
-For each shell thickness, the absorption is suppressed at small input field intensities and enhanced at high intensities, when the collective emission and the absorber-to-particle energy transfer are suppressed due to the absorber transition saturation.
-Let us see what happens with the decrease of the absorber state de-population rate (i.e. the increase $\tau$).
-Figure 5: $n_{1}$ (a) and $R$ (b) versus the normalized input field intensity for $N_{m}=10$ and different absorber upper level de-population time $\tau=2$ ps (curves 1), $20$ ps (2) and 50 ps (3).
+Figure 4 shows the dependence of $n_{1}$ and $R$ on the normalised external field intensity $\sigma_{m}P\tau$ for a large number of absorbers $N_{m}=120$ and different nanoparticle shell thicknesses $\Delta r=5$ (curves 1), $6$ (2) and $9$ nm (3).
+We see that the bistability disappears with only 1 nm increase in the nanoparticle shell thickness.
+For each shell thickness, the absorption is suppressed at low input field intensities and enhanced at high intensities, when the collective emission and the absorber-to-particle energy transfer are suppressed due to the absorber transition saturation.
+Let us see what happens with the decrease of the depopulation rate of the absorber state (i.e. the increase of $\tau$).
+Figure 5: $n_{1}$ (a) and $R$ (b) versus the normalised input field intensity for $N_{m}=10$ and different absorber upper level depopulation time $\tau=2$ ps (curves 1), $20$ ps (2) and 50 ps (3).
 Curve 4 is for an absorber without the nanoparticle.
-The shell thickness $\Delta r=5$ nm, the rest of parameters is the same as for Figure 3.
-The upper-level population and the absorption are reduced with the increase of $\tau$.
-The absorption enhancement by LPR is reduced with the increase of $\tau$, from curves 1 to curves 3 in Fig 5, because of the growing of the spontaneous emission to the LPR contribution to the population decay.
-Suppose we have a large absorber upper state de-population time $\tau=1/2\gamma_{sp}=0.5$ ns; such a de-population is, for example, because of the spontaneous emission to free space with the typical optical domain rate $\gamma_{sp}=1$ GHz Scully and Zubairy (1997).
-We assume the absorber transition half-width $\gamma=100$ GHz.
+The shell thickness $\Delta r=5$ nm, the rest of the parameters are the same as in Figure 3.
+The upper level population and the absorption are reduced with the increase of $\tau$.
+We first consider the case without the bistability and with a large absorption enhancement, as shown in Fig.
+3 by the curve 1 for $N_{m}=10$ and $\tau=2$ ps.
+The absorption enhancement by LPR is decreases with increasing $\tau$, from curves 1 to curves 3 in Fig 5, due to the increase of the spontaneous emission to the LPR contribution to the population decay.
+Let us now increase the absorber state depopulation time $\tau$ at a large number of absorbers $N_{m}=120$, when the condition (32) is satisfied and we have the absorption bistability.
+We see in Fig.
+6 that the bistability region grows with the increase of $\tau$, while the absorption enhancement is very small.
+Figure 6: The upper absorber state population $n_{1}(\sigma_{m}P\tau)$ at the absorption bistability, when $N_{m}=120$ and different $\tau=2$ ps (curve 1), $20$ ps (curve 2) and $50$ ps (curve 3).
+Unstable parts of the curves are indicated by the dashed lines.
+The bistability region increases as $\tau$ increases.
+Curve 4 is for the absorber in free space.
+The absorption enhancement (i.e.
+the positive difference in the upper parts of curves 1-3 and 4) is quite small.
+Suppose we have a large absorber upper state depopulation time $\tau=1/2\gamma_{sp}=0.5$ ns; such depopulation occurs, for example, due to spontaneous emission into free space with the typical optical domain rate $\gamma_{sp}=1$ GHz [52].
+We assume that the absorber transition has a half-width $\gamma=100$ GHz.
 Then $\alpha_{p}/\alpha_{m}=20<2\gamma\tau=100$ so the condition (29) is not satisfied and there is no absorption enhancement by the LPR at any external field intensity.
-Meanwhile, the bistability in absorption exists, as shown by curve 3 in Figure 6, when the bistability condition (32) is true.
-Figure 6: The upper absorber state population $n_{1}(\sigma_{m}P\tau)$ at the suppression of absorption by the spontaneous emission to LPR mode, when $\tau=0.5$ ns and $\gamma=100$ GHz.
+Meanwhile, there is the bistability in the absorption, as shown by curve 3 in Figure 7, when the bistability condition (32) is satisfied.
+Figure 7: The upper absorber state population $n_{1}(\sigma_{m}P\tau)$ with suppression of absorption by the spontaneous emission in the LPR mode, when $\tau=0.5$ ns and $\gamma=100$ GHz.
 The number of absorbers $N_{m}=10$ is for curve 1; $N_{m}=16$ is for curve 2 and $N_{m}=30$ is for curve 3; curve 4 is for the absorber in free space.
-The bistability condition (32) is satisfied and the bistability in absorption exists for curve 3.
-The dashed part of curve 3 corresponds to the unstable stationary state.
-VI Discussion
+The bistability condition (32) is satisfied and there is bistability in absorption exists for curve 3.
+The dashed part of curve 3 corresponds to the unstable steady state.
+There is no the absorption enhancement.
+6 Discussion
 An electric field enhancement near plasmonic nanoparticles increases the resonant absorption of atoms, molecules, semiconductor media, etc.
-The estimation (27) shows the growth of the absorption cross-section near plasmonic particle in a factor $\alpha_{p}^{2}/4r^{6}=9.5$ – without taking into account the spontaneous emission of the absorber to the LPR mode (described by the denominator in Eq.
+The estimation (27) shows the increase of the absorption cross section near plasmonic particle by a factor $\alpha_{p}^{2}/4r^{6}=9.5$ – without taking into account the spontaneous emission of the absorber to the LPR mode (described by the denominator in Eq.
 (27)).
-Spontaneous and collective emissions to the LPR mode significantly reduce the absorption making it, at some conditions, smaller than in the free space.
-In our examples, the maximum absorption enhancement near plasmonic nanoparticles is about the factor 3 (see curve 1 in Figure 3b), with the spontaneous emission to the LPR mode taken into account.
+Spontaneous and collective emissions to the LPR mode significantly reduce the absorption making it smaller than in the free space under some conditions.
+In our examples, the maximum absorption enhancement near plasmonic nanoparticles is about a factor of 3 (see curve 1 in Figure 3b), taking into account the spontaneous emission to the LPR mode.
 The maximum absorption enhancement is found for a small external field when the absorbing transition is not saturated.
-A noticeable absorption enhancement is possible at a fast (on ps time scale) de-population of the upper absorbing state, when the rate of spontaneous emission to the LPR mode is about or less than the de-population rate due to other processes as, for example, the decay of absorber excitation to the non-resonant high-order plasmonic modes Dulkeith et al.
-(2005).
+A noticeable absorption enhancement is possible with a fast (on ps time scale) depopulation of the upper absorbing state, when the rate of spontaneous emission to the LPR mode is close to or less than the depopulation rate due to other processes.
 When the population decay is not so fast, the relative suppression of the absorber excitation by the radiation to the LPR is large and we do not have the absorption enhancement, as shown in Figure 7.
-The condition (29) must be satisfied for the LPR absorption enhancement.
-Spontaneous and collective emissions to the LPR mode are important factors, they can not be neglected in the resonant absorption near metal nanoparticles.
-The spontaneous emission enhancement by the cavity mode is well-known in the laser theory Oraevskii (1994) as a Pursell effect Purcell (1946).
-The metal nanoparticle is analogous to the cavity for the LPR mode, so such a cavity enhances the spontaneous emission similar to a laser cavity.
-When a few resonant absorbers are placed near metal nanoparticles, a collective emission from absorbers to the plasmonic mode is possible.
-A photon emitted by the atom $i$ is re-absorbed by the atom $j$ and emitted again by the collective emission.
-This is a Dicke effect Gross and Haroche (1982) with the probability (and, therefore, the rate) larger than for the photon emission by a single atom.
-The collective emission de-populates the absorber’s upper state.
-Collective emission to the LPR mode is similar to the collective emission to a lasing mode Leymann et al.
-(2015); Jahnke et al.
-(2016); André et al.
-(2019); Protsenko and Uskov (2022).
-The collective emission is a nonlinear process, its rate depends on the emitting state populations.
-In particular, the collective emission is suppressed at the saturation of the emitting transition.
-The particle-absorber near-field interaction is a mutual process: the metal nano-particle LPR mode excites the polarization and transfers the energy to the absorber through the near field and vice-versa: absorbers excite the particle polarisation through the near field.
+For LPR absorption enhancement, condition (29) must be satisfied.
+Spontaneous and collective emissions to the LPR mode are important factors that cannot be neglected in the resonant absorption near metal nanoparticles.
+The enhancement of spontaneous emission by the cavity mode is well known in the laser theory [59] as the Pursell effect [60].
+The metal nanoparticle is analogous to the cavity for the LPR mode, so such a cavity will enhance the spontaneous emission similar to a laser cavity.
+When a few resonant absorbers are placed close to metal nanoparticles, collective emission from the absorbers into the plasmonic mode is possible.
+A photon emitted by the atom $i$ is re-absorbed by the atom $j$ and re-emitted by the collective emission.
+This is a Dicke effect [61] with the probability (and hence the rate) is greater than for the photon emission from a single atom.
+The collective emission depopulates the upper state of the absorber.
+Collective emission to the LPR mode is similar to the collective emission to a lasing mode [62, 63, 28, 29].
+Collective emission is a non-linear process, its rate depends on the populations of the emitting states.
+In particular, collective emission is suppressed at the saturation of the emitting transition.
+The particle-absorber near-field interaction is a two-way process: the LPR mode of the metal nano-particle excites the polarisation and transfers the energy through the near-field of the absorber, and vice versa: absorbers excite the particle polarisation through the near-field.
 Such an energy transfer is the origin of the suppression factor in the denominator in Eq.
 (14) for the absorber polarisation.
-The collective emission and the emitter-particle interaction depend on the absorber states population.
-Such a dependence leads to a positive feedback: the absorption is increased with the upper-level population at certain conditions.
-The feedback is fast: note the square in the nonlinear denominator in Eq.
+The collective emission and the absorber-particle interaction depend on the absorber state population.
+Such a dependence leads to a positive feedback: the absorption is increased with the upper level population under certain conditions.
+The feedback is fast: note the square in the non-linear denominator in Eq.
 (26).
-So, at certain parameter values, the feedback nonlinearity "overtakes" the transition saturation and leads to a decrease of the upper-level population $n_{1}$ with the external field intensity.
-When $n_{1}$ decreases, the stationary absorption is unstable, see the dashed parts of curves in Figs.
+Thus, for certain parameter values, the feedback nonlinearity ”overtakes” the transition saturation and leads to a decrease of the upper-level population $n_{1}$ with the external field intensity.
+As $n_{1}$ decreases, the steady-state absorption is unstable, see the dashed parts of the curves in Figs.
 3, 4, and 6.
 Such anomalous absorption causes absorption bistability.
-The bistability exists at the condition (32) which can be satisfied at the enhancement (when condition (29) is true) or the suppression (when condition (29) is not true) of absorption by LPR, see Figures 4 and 6.
-With the LPR absorption enhancement, the bistability is between the absorption suppression and the absorption enhancement regimes, see Figures 3 and 4.
-The optical bistability is possible in plasmonic nanolaser with external signal Protsenko et al.
-(2008).
-The model of Protsenko et al.
-(2008) involves a single emitter near the plasmonic nanoparticle and does not include the spontaneous and collective emissions to the LPR mode.
-The physical nature of a well-known absorptive optical bistability in a cavity with a two-level absorber (or absorbers) Koch, Schmidt, and Haug (1985) is the resonant transition saturation.
-This is different from the physics of the bistability in the resonant absorption in the LPR near metal nanoparticle described here.
-The absorptive bistability Koch, Schmidt, and Haug (1985) is not related to the collective emission and the near-field interaction of absorbers with the metal nanoparticle.
-The nanoparticle with resonant emitters is, probably, one of the smallest systems with optical bistability with the size much smaller than the optical wavelength.
-Here we carry out the semi-classical analysis of the resonant absorption near a metal nanoparticle.
-Only the term of the spontaneous emission into the LPR mode in Eq.
-(17) is introduced according to quantum mechanical relation (16).
-We note that the term of the spontaneous emission to the lasing mode appears in the laser theory the same way P.
-R.
-Rice  and Carmichael (1994); André et al.
-(2019).
-More detailed theory of the resonant absorption near plasmonic nanoparticles must include quantum fluctuations of the field and the absorber states populations.
-Such a quantum analysis is necessary, taking into account a small volume of the plasmonic mode Maier (2006), which keeps only one or a few photons.
-Such an analysis will be carried out in the future.
-VII Conclusion
-We investigate the enhancement of the resonant absorption of the electromagnetic field by two-level absorbers (atoms or molecules) near plasmonic nanoparticles, taking into account the spontaneous and collective emission from absorbers to the localized plasmon resonance (LPR) mode of the particle.
-We derive quantum equations for the particle and the resonant absorber polarisation and the absorbing state populations.
-We analyze the structure of equations and describe the physical mechanisms of the stationary absorption and the particle-absorber interaction.
-We find that the absorption enhancement by the LPR is possible under certain conditions.
-The absorption suppression is possible as well due to the spontaneous, collective emissions of absorbers to the LPR mode and the energy transfer from absorbers to the particle.
-We find conditions necessary for the absorption enhancement.
-One condition is a fast, of the pico-second time scale, de-population of the excited absorber states, for example, by some non-radiative processes.
-We find conditions and investigate the bistability in the resonant absorption near the metal nanoparticle.
-The bistability is caused by the collective emission and the nonlinear interaction between the particle and absorbers through the LPR near field.
-The bistability appears at a large number of absorbers near the nanoparticle and possible with the absorption enhancement or suppression.
-Results can be used everywhere, when the resonant absorption near metal nanoparticles is used, for example, for the plasmon-stimulated photo-catalysis Zhang et al.
-(2013) or in the plasmonic photo-voltaic Jang et al.
-(2016).
-Acknowledgements.
+The bistability exists under the condition (32), which can be satisfied by the enhancement (when condition (29) is true) or the suppression (when condition (29) is not true) of absorption by LPR, see Figures 4 and 6.
+In the case of LPR absorption enhancement, the bistability is between the absorption suppression and absorption enhancement regimes, see Figures 3 and 4.
+Optical bistability is possible in plasmonic nanolasers with an external signal [26].
+The model of [26] includes a single emitter near the plasmonic nanoparticle and does not consider the spontaneous and collective emissions to the LPR mode.
+The physical nature of the well-known absorptive optical bistability in a cavity with a two-level absorber (or absorbers) [64] is the saturation of the resonant transition.
+This is different from the physics of the resonant absorption bistability in the LPR near metal nanoparticles described here.
+The absorptive bistability [64] is not related to the collective emission and the near-field interaction of the absorbers with the metal nanoparticle.
+The nanoparticle with the nearby resonant absorbers (molecules) is probably one of the smallest systems with optical bistability, with a size much smaller than the optical wavelength.
+We have shown that the absorption enhancement near the metal nanoparticle and the bistability are strongly influenced by nonlinear collective absorption of the nanoparticle and nearby molecules, which occurs at high light power densities ($>1~{}MW/cm^{2}$).
+Such power densities for single spherical metal nanoparticles can only be achieved with powerful pulsed lasers.
+This makes potential applications in photovoltaics, photocatalysis, sensors, etc difficult.
+At the same time, the necessary power densities are easily achieved in plasmonic nanosystems with strong spatial localisation of the electromagnetic field in a plasmonic gap mode even at moderate levels of external light excitation ($\sim 1~{}W/cm^{2}$ and less) [16, 17, 18, 19].
+This makes the study of the collective absorption near the plasmonic gap mode to be an interesting and important task.
+The present approach can be directly generalised to the plasmonic gap case.
+Here we perform a semi-classical analysis of the resonance absorption near a metal nanoparticle.
+Only the term of the spontaneous emission in the LPR mode in Eq.
+(17) is introduced according to the quantum mechanical relation (16).
+We note that the term of the spontaneous emission into the laser mode appears in the same way in the laser theory [65, 28].
+A more detailed theory of the resonant absorption near plasmonic nanoparticles must include quantum fluctuations of the field and the absorber state populations.
+Such a quantum analysis is necessary, taking into account a small volume of the plasmonic mode [66], holding only one or a few photons.
+Such an analysis will be performed in the future.
+7 Conclusion
+We study the enhancement of the resonant absorption of the electromagnetic field by two-level absorbers (atoms or molecules) in the vicinity of plasmonic nanoparticles, taking into account the spontaneous and collective emission from the absorbers to the localised plasmon resonance (LPR) mode of the particle.
+We derive quantum equations for the polarisation of the particle and the resonant absorber, and for the populations of absorbing states.
+We analyse the structure of the equations and describe the physical mechanisms of the stationary absorption and the particle-absorber interaction.
+We find that absorption enhancement by LPR is possible under certain conditions.
+Absorption suppression is also possible due to individual spontaneous and collective emissions from absorbers to the LPR mode and energy transfer from absorbers to the particle.
+We find conditions necessary for absorption enhancement.
+One condition is a fast, on the picosecond time scale, depopulation of the excited absorber states, e.g.
+by some non-radiative processes.
+We find the conditions and study the bistability in the resonant absorption near the metal nanoparticle.
+The bistability is caused by the collective emission and the nonlinear interaction between the particle and the absorbers through the LPR near field.
+The bistability occurs with a strong external field excitation, a large number of absorbers in the vicinity of the nanoparticle, and is possible with absorption enhancement or suppression.
+The results can be applied wherever resonant absorption near metal nanoparticles is used, for example in plasmon-stimulated photocatalysis [8] or plasmonic photovoltaics [7].
+Acknowledgments
 I.E.P., A.V.U. and N.V.N. acknowledge the Russian Science Foundation (Grant No.20-19-00559) for the support.
 Appendix. Nanoparticle resonant polarisability
-Polarizability $\alpha_{p}$ of the spherical nanoparticle of the radius $r_{p}$ in the electric field with the wavelength $\lambda$, in the dielectric environment with the refractive index $\eta$ is Protsenko and Uskov (2012)
+The polarizability $\alpha_{p}$ of a spherical nanoparticle of the radius $r_{p}$ in an electric field of wavelength $\lambda$, in a dielectric medium of refractive index $\eta$ is [35]
 $\alpha_{p}(\lambda,r_{p})=r_{p}^{3}\frac{\varepsilon_{p}(\lambda,r_{p})/\eta^{2}-1}{[1-iR_{rad}(\lambda,r_{p})][\varepsilon_{p}(\lambda,r_{p})/\eta^{2}+2]}$ (34)
-where $\varepsilon_{p}$ is the dielectric function of the material of the nanoparticle and the factor
+where $\varepsilon_{p}$ is the dielectric function of the nanoparticle material and the factor
 $R_{rad}(\lambda,r_{p})=\frac{16\pi^{3}}{3}\left(\frac{\eta r_{p}}{\lambda}\right)^{3}\frac{\varepsilon_{p}(\lambda,r_{p})/\eta^{2}-1}{\varepsilon_{p}(\lambda,r_{p})/\eta^{2}+2}$ (35)
 describe the radiation losses.
-The nanoparticle material dielectric function $\varepsilon_{p}(\lambda,r_{p})$ takes into account the losses due to collisions of the nanoparticle electrons with the nanoparticle boundary Protsenko and Uskov (2012); Uskov et al.
-(2014)
+The nanoparticle material dielectric function $\varepsilon_{p}(\lambda,r_{p})$ taking into account the losses due to collisions of the nanoparticle electrons with the nanoparticle boundary [35, 36] is
 $\varepsilon_{p}(\lambda,r_{p})=\varepsilon_{0}-\frac{(\lambda/\lambda_{p})^{2}}{(1+i\lambda/\lambda_{c})(r_{c}/r_{p}+1)}.$ (36)
-[…]
+The expression (36) with $r_{c}=0$ is a Drude dielectric function.
+For silver $\varepsilon_{0}=5$, $\lambda_{p}=0.14$ mkm, $\lambda_{c}=32$ mkm [67].
+The term $\sim r_{c}=A_{0}v_{F}\lambda_{c}/2\pi c$ describes the correction to the dielectric function due to collisions of electrons with the nanoparticle surface.
+The factor $A_{0}=0.7$, $v_{F}\approx 1.3\cdot 10^{6}$ m/s is the Fermi velocity of the electron, $c$ is the speed of light in vacuum [35].
+Radiation losses increase with the nanoparticle radius $r_{p}$ while the losses due to collisions with the nanoparticle surface decrease with $r_{p}$.
+This, there is an optimum $r_{p}$, when the nanoparticle LPR has the maximum quality factor $Q_{LPR}$.
+Estimates show that the maximum $Q_{LPR}\approx 29$, when $r_{p}=r_{p}^{opt}=12$ nm.
+The LPR wavelength $\lambda_{LPR}\approx 408$ nm and the LPR half-width $\Gamma\approx 80$ THz.
+The absolute value of the maximum resonant polarisability of the nanoparticles is $\alpha_{p}=|\alpha_{p}(\lambda_{LPR},r_{p}^{opt})|\approx 3\cdot 10^{-5}$ mkm3.